### v2

# MIBench: A Comprehensive Benchmark for Model Inversion Attack and Defense

Abstract
Model Inversion (MI) attacks aim at leveraging the output information of target models to reconstruct privacy-sensitive training data, raising widespread concerns on privacy threats of Deep Neural Networks (DNNs). Unfortunately, in tandem with the rapid evolution of MI attacks, the lack of a comprehensive, aligned, and reliable benchmark has emerged as a formidable challenge. This deficiency leads to inadequate comparisons between different attack methods and inconsistent experimental setups. In this paper, we introduce the first practical benchmark for model inversion attacks and defenses to address this critical gap, which is named MIBench. This benchmark serves as an extensible and reproducible modular-based toolbox and currently integrates a total of 16 state-of-the-art attack and defense methods. Moreover, we furnish a suite of assessment tools encompassing 9 commonly used evaluation protocols to facilitate standardized and fair evaluation and analysis. Capitalizing on this foundation, we conduct extensive experiments from multiple perspectives to holistically compare and analyze the performance of various methods across different scenarios, which overcomes the misalignment issues and discrepancy prevalent in previous works. Based on the collected attack methods and defense strategies, we analyze the impact of target resolution, defense robustness, model predictive power, model architectures, transferability and loss function. Our hope is that this MIBench could provide a unified, practical and extensible toolbox and is widely utilized by researchers in the field to rigorously test and compare their novel methods, ensuring equitable evaluations and thereby propelling further advancements in the future development.
1 Introduction
In recent years, Model Inversion (MI) attacks have raised alarms over the potential privacy breaches of sensitive personal information, including the leakage of privacy images in face recognition models (He et al., 2016), sensitive health details in medical data (Wang et al., 2022), financial information such as transaction records and account balances (Ozbayoglu et al., 2020), and personal preferences and social connections in social media data (Feng et al., 2022). In the MI attacks, an attacker aims to infer private training data from the output information of the target model. Fredrikson et al. (2014) proposed the first MI attack against linear regression models to reconstruct sensitive features of genomic data. Subsequent studies (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) have extended MI attacks to more Machine Learning (ML) models. Zhang et al. (2020b) first introduces the GANs as stronger image priors, paving the way for more applications in GAN-based methods (Chen et al., 2021a; Yuan et al., 2022; 2023b; Wang et al., 2021b).
However, some critical challenges are posed owing to the lack of a comprehensive, fair, and reliable benchmark. Specifically, the evaluation of new methods is often confined to comparisons with a narrow selection of prior works, limiting the scope and depth of analysis. For instance, some methods (Nguyen et al., 2023b; Yuan et al., 2023a) exhibit superior performance for lower-resolution images while other methods (Struppek et al., 2022; Qiu et al., 2024) perform better at higher resolutions. However, these studies only evaluate under their predominant resolutions, and thus do not provide a unified and comprehensive comparison. Additionally, the absence of unified experimental protocols results in a fragmented landscape where there is less validity and fairness in the comparative studies. For example, the original GMI (Zhang et al., 2020a) achieves the attack accuracy of 28%, 44%, 46% when attacking three different classifiers trained on the CelebFaces Attributes(CelebA) (Liu et al., 2015), while merely maintaining lower attack accuracy of 21%, 32%, 31% and 21%, 31%, 29% when compared in the KEDMI (Chen et al., 2021b) and PLGMI (Yuan et al., 2023a) under the same experimental setup. Compounding these issues, discrepancies in the evaluation metrics further obscure the reliability of reported conclusions, potentially steering the field towards biased or misleading insights. These shortcomings impede both the accurate measurement of advancements in the MI field and the systematic exploration of its theoretical underpinnings, underscoring the urgent need for a harmonized framework to facilitate robust and transparent research practices.
To alleviate these problems, we establish the first benchmark in MI field, named MIBench. For building an extensible modular-based toolbox, we disassemble the pipeline of MI attacks and defenses into four main modules, each designated for data preprocessing, attack methods, defense strategies and evaluation, hence enhancing the extensibility of this unified framework. The proposed MIBench has encompassed a total of 16 distinct attack and defense methods, coupled with 9 prevalent evaluation protocols to adequately measure the comprehensive performance of individual MI methods. Furthermore, we conduct extensive experiments from diverse perspectives to achieve a thorough appraisal of the competence of existing MI methods, while simultaneously venturing into undiscovered insights to inspire potential avenues for future research. We expect that this reproducible benchmark will facilitate the further development of MI field and bring more innovative explorations in the subsequent study. Our main contributions are as follows:
We build the first comprehensive benchmark in MI field, which serves as an extensible and reproducible modular-based toolbox for researchers. We expect that this will foster the development of more powerful MI attacks by making it easier to evaluate their effectiveness across multiple distinct dimensions.
We implement 16 state-of-the-art attack methods and defense strategies and 9 evaluation protocols currently in our benchmark. We hope that the benchmark can further identify the most successful ideas in defending against rapid development of potential MI attacks.
We conduct extensive experiments to thoroughly assess different MI methods under multiple settings and study the effects of different factors to offer new insights on the MI field. In particular, we validate that stronger model predictive power correlates with an increased likelihood of privacy leakage. Moreover, our analysis reveals that certain defense algorithms also fail when the target model achieves high prediction accuracy.
2 Related Work
Model Inversion Attacks. In the MI attacks, the malicious adversary aims to reconstruct privacy-sensitive data by leveraging the output prediction confidence of the target classifier and other auxiliary priors. Normally, the attacker requires a public dataset that shares structural similarities with the private dataset but without intersecting classes to pre-train the generator. For example, an open-source face dataset serves as essential public data when targeting a face recognition model. For a typical GAN-based MI attack, attackers attempt to recover private images from the GAN’s latent vectors initialized by Gaussian distribution, given the target image classifier parameterized with weights and the trained generator . The attack process can be formulated as follows:
where is the target class, typically denotes the classification loss, is a hyperparameter, and is the prior knowledge regularization (e.g., the discriminator’s classification loss) used to improve the reality of . By minimizing the above equation, the adversary updates the latent vectors into the optimal results and generate final images through .
Fredrikson et al. (2014) first introduce the concept of MI attacks in the context of genomic privacy. They find that maximizing the posterior probabilities of a linear regression model can reconstruct the original genomic markers. Ensuing works (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) manage to design MI attacks for more kinds of models and private data, but are still limited to attacking simple networks and grayscale images. To enhance the reconstruction performance on complex RGB images, GMI (Zhang et al., 2020a) first propose to incorporate the rich prior knowledge (Fang et al., 2023; Gu et al., 2020; Fang et al., 2024) within the pre-trained Generative Adversarial Networks (GANs) (Goodfellow et al., 2014). Specifically, GMI starts by generating a series of preliminary fake images, and then iteratively optimizes the input latent vectors that are used for generation. Based on GMI, KEDMI (Chen et al., 2021b) refine the discriminator by introducing target labels to recover the distribution of the input latent vectors. VMI (Wang et al., 2021a) utilize variational inference to model MI attacks and adopts KL-divergence as the regularization to better estimate the target distribution. PPA (Struppek et al., 2022) introduce a series of techniques such as initial selection, post-selection, and data argumentation to enhance MI attacks and manages to recover high-resolution images by the pre-trained StyleGAN2 (Karras et al., 2019). LOMMA (Nguyen et al., 2023b) integrate model augmentation and model distillation into MI attacks to tackle the problem of over-fitting. PLGMI (Yuan et al., 2023a) leverage a top- selection technique to generate pseudo labels to further guide the training process of GAN.
Besides, based on whether the parameters and structures of the victim model are fully accessible to the attackers, MI attacks can be further split into white-box attacks and black-box attacks. Note that in black-box settings, the gradients can no longer be computed by the back-propagation process. Thus, Yuan et al. (2022) address this problem by sampling numerous latent vectors from random noise, selecting the ones that can generate correct labels, and updating the latent vectors solely on discriminator loss. Nguyen et al. (2023a) propose to conduct MI attacks on various surrogate models instead of the unfamiliar victim model, transforming the black-box settings into white-box settings.
Model Inversion Defenses. To defend MI attacks, most existing methods can be categorized into two types: model output processing (Yang et al., 2020; Wen et al., 2021; Ye et al., 2022) and robust model training (Gong et al., 2023; Titcombe et al., 2021; Li et al., 2022; Wang et al., 2021c; Peng et al., 2022; Struppek et al., 2023). Model output processing refers to reducing the private information carried in the victim model’s output to promote privacy. Yang et al. (2020) propose to train an autoencoder to purify the output vector by decreasing its degree of dispersion. Wen et al. (2021) apply adversarial noises to the model output and confuse the attackers. Ye et al. (2022) leverage a differential privacy mechanism to divide the output vector into multiple sub-ranges. Robust model training refers to that incorporating the defense strategies during the training process. MID Wang et al. (2021c) penalizes the mutual information between model inputs and outputs in the training loss, thus reducing the redundant information carried in the model output that may be abused by the attackers. However, simply decreasing the dependency between the inputs and outputs also results in model performance degradation. To alleviate this issue and strike a better balance between model utility and user privacy, Bilateral Dependency Optimization (BiDO) (Peng et al., 2022) minimizes the dependency between the inputs and outputs while maximizing the dependency between the latent representations and outputs. Gong et al. (2023) propose to leverage GAN to generate fake public samples to mislead the attackers. Titcombe et al. (2021) defend MI attacks by adding Laplacian noise to intermediate representations. LS (Struppek et al., 2023) finds that label smoothing with negative factors can help privacy preservation. TL (Ho et al., 2024) leverages transfer learning to limit the number of layers encoding sensitive information and thus improves the robustness to MI attacks.
3 Our Benchmark
3.1 Dataset
Considering existing MI attacks primarily focus on reconstructing private facial data from image classifiers, we select 4 widely recognized face datasets as the basic datasets for our benchmark, which include Flickr-Faces-HQ (FFHQ) (Karras et al., 2019), MetFaces (Karras et al., 2020a), FaceScrub (Ng & Winkler, 2014), and CelebFaces Attributes (CelebA) (Liu et al., 2015). Generally, FFHQ and MetFaces are employed as public datasets for pre-training auxiliary priors, whereas FaceScrub and CelebA serve as the target private datasets to attack. Our benchmark facilitates researchers to freely combine public datasets with private ones, thereby enabling customized experimental setups. Extensive evaluation on more non-facial datasets are presented in Sec.C.3
Notably, the target resolutions across different MI attacks are not uniform. The majority of attack methods concentrate on low-resolution images of , while recent attack methods have begun to focus on higher resolutions, such as . Therefore, our benchmark offers 2 versions of low-resolution and higher-resolution for the aforementioned 4 datasets and prepares multiple transformation tools for processing images, freeing researchers from the laborious tasks of data preprocessing. More details regarding the datasets can be found in the Sec. B.2 of the Appendix.
3.2 Implemented Methods
Our benchmark includes a total of 16 methods, comprising 11 attack methods and 4 defense strategies. With a focus on Generative Adversarial Network (GAN)-based MI attacks, we selectively reproduce methods from recent years that have been published in top-tier conferences or journals in the computer vision or machine learning domains. This criterion ensures the reliability and validity of the implemented methods. Considering the main targets in our benchmark are image classifiers for RGB images, the learning-based MI attacks (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) are not incorporated currently. More detailed information about the implemented methods is stated in Sec. A.2 and Sec. A.3.
Attacks. Based on the accessibility to the target model’s parameters, we categorize MI attacks into white-box and black-box attacks. White-box attacks can entail full knowledge of the target model, enabling the computation of gradients for performing backpropagation, while black-box attacks are constrained to merely obtaining the prediction confidence vectors of the target model. Our benchmark includes 8 white-box attack methods and 4 black-box attack methods, as summarized in Table 1.
| Attack Method | Accessibility | Reference | GAN Prior | Official Resolution |
| GMI (Zhang et al., 2020a) | White-box | CVPR-2020 | WGAN-GP | |
| KEDMI (Chen et al., 2021b) | White-box | ICCV-2021 | Inversion-specific GAN† | |
| VMI (Wang et al., 2021a) | White-box | NeurIPS-2021 | StyleGAN2 | |
| Mirror∗ (An et al., 2022) | White-box/Black-box | NDSS-2022 | StyleGAN | |
| PPA (Struppek et al., 2022) | White-box | ICML-2022 | StyleGAN2 | |
| PLGMI (Yuan et al., 2023a) | White-box | AAAI-2023 | Conditional GAN | |
| LOMMA (Nguyen et al., 2023b) | White-box | CVPR-2023 | ||
| IF-GMI (Qiu et al., 2024) | White-box | ECCV-2024 | StyleGAN2 | |
| BREPMI (Kahla et al., 2022) | Black-box | CVPR-2022 | WGAN-GP | |
| C2FMI (Ye et al., 2023) | Black-box | TDSC-2023 | StyleGAN2 | |
| RLBMI (Han et al., 2023) | Black-box | CVPR-2023 | WGAN-GP | |
| LOKT (Nguyen et al., 2023a) | Black-box | NeurIPS-2023 | ACGAN |
Defenses. To effectively defend against MI attacks, the defender typically employs defense strategies during the training process of victim classifiers. Our benchmark includes 4 typical defense strategies and the details are presented in Table 2.
| Defense Strategy | Reference | Core Technique | Description |
| MID (Wang et al., 2021c) | AAAI-2021 | Regularization | Utilize mutual information regularization to limit leaked information about the model input in the model output |
| BiDO (Peng et al., 2022) | KDD-2022 | Regularization | Minimize dependency between latent vectors and the model input while maximizing dependency between latent vectors and the outputs |
| LS (Struppek et al., 2023) | ICLR-2024 | Label Smoothing | Adjusting the label smoothing with negative factors contributes to increasing privacy protection |
| TL (Ho et al., 2024) | CVPR-2024 | Transfer Learning | Utilize transfer learning to limit the number of layers encoding privacy-sensitive information for robustness to MI attacks |
3.3 Toolbox
We implement an extensible and reproducible modular-based toolbox for our benchmark, as shown in Fig 1. The framework can be divided into four main modules, including Data Preprocess Module, Attack Module, Defense Module and Evaluation Module.
Data Preprocess Module. This module is designed to preprocess all data resources required before launching attacks or defenses, including datasets, classifiers and parameters. Consequently, we furnish this module with three fundamental functionalities: dataset preprocessing, target model training, and parameter management. For dataset preprocessing, we build a unified pipeline for each dataset, which automatically carries out a series of operations such as spliting dataset and image transformations (e.g. center crop and resize) based on the split file and chosen resolution from users. For target model training, users can further leverage the processed datasets to train designated classifiers. We abstract the general procedures in classifier training into a base trainer class to facilitate users in extending and customizing their own classifiers. For parameter management, we encapsulate parameters used in different processes into specific configuration classes, such as TrainConfig designated for the training process and AttackConfig for the attack process, thus ensuring organized and efficient parameter handling across various operations in the workflow.
Attack Module. The workflow of MI attacks can be roughly divided into two stages. The first stage is GAN training, where the module abstracts the general training process of GANs into a basic trainer class. This allows users not only train GANs that are pre-built into the benchmark, but also extend to their uniquely designed GANs. The second stage is the core inversion attack, which we split into three parts: latent vectors initialization, iterative optimization, and an optional post-processing step. After completion of the attack, the module preserves essential data such as the final optimized images and latent vectors, facilitating subsequent evaluation and analysis.
Defense Module. Considering the mainstream MI defense strategies are applied during the training process of target classifiers, we design the defense module following the target model training functionality within the Data Preprocessing Module. To enhance extensibility, we incorporate defense strategies as part of the training parameters for classifiers to enable the defense during the training process of target models, which decouples the defense from the training pipeline. In this way, we allow users to customize their own defense strategies against MI attacks.
Evaluation Module. Our benchmark concentrates on the evaluation at distribution level instead of sample level, assessing the overall performance of the whole reconstructed dataset. Therefore, we provide a total of 9 widely recognized distribution level evaluation metrics for users, which can be categorized into four types according to the evaluated content: Accuracy, Feature Distance, FID (Heusel et al., 2017) and Sample Diversity. For Accuracy, this metric measures how well the reconstructed samples resemble the target class, consisting of and . For Feature Distance, it computes the average shortest distance from features between each reconstructed sample and private data to measure the similarity between the feature space of reconstructed samples and private dataset. For FID, the lower FID score shows higher realism and overall diversity (Wang et al., 2021a). For Sample Diversity, the higher values indicate greater intra-class diversity. Moreover, we provide convenient tools for analysis, including Standard Deviation Calculation and Visualization.
4 Experiment
4.1 Experimental Setups
To ensure fair and uniform comparison and evaluation, we select the FFHQ (Karras et al., 2019) as the public dataset and FaceScrub (Ng & Winkler, 2014) as the private dataset for all the experiments in the Experiment section. The target models are fixed to the IR-152 (He et al., 2016) for low-resolution scenario and ResNet-152 (He et al., 2016) for high-resolution scenario, both trained on the FaceScrub. For each attack method, the number of images reconstructed per class is set to 5 due to considerations of time and computation cost. More detailed experimental settings are listed in Sec.B in the Appendix.
Notably, we limit the evaluation exhibited in the Experiment section to merely three metrics, including Accuracy, Feature Distance and FID (Heusel et al., 2017), while Sample Diversity is presented in the Sec.C.1 in the Appendix. Moreover, the VMI (Wang et al., 2021a) and RLBMI (Han et al., 2023) will be further evaluated in Sec. C.7 owing to their excessive need of time.
4.2 Evaluation on Different Attack Methods
In this part, we prepare a unified experimental setting for different MI attack methods to conduct a fair comparison. The resolution of private and public datasets is set to , indicating a relatively easier scenario. Comparisons of white-box and black-box MI attacks are presented in Table 3.
Remarkably, the PLGMI (Yuan et al., 2023a) and LOKT (Nguyen et al., 2023a) achieve state-of-the-art comprehensive performance in white-box attacks and black-box attacks respectively, showing significant superiority in Accuracy and Feature Distance metrics. However, the lowest FID scores occur in the PPA (Struppek et al., 2022) and C2FMI (Ye et al., 2023) respectively instead of the above methods. We infer that this is because PPA and C2FMI employ more powerful generators (e.g. StyleGAN2 (Karras et al., 2020b)) as the GAN prior compared to PLGMI and LOKT, leading to more real image generation. Visualization in Fig 2 further validates the inference.
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror(white) | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI | |||||
| BREPMI | |||||
| Mirror(black) | |||||
| C2FMI | 43.488 | ||||
| LOKT |
4.3 Evaluation on Higher Resolution
Recent attack methods have attempted to conquer higher resolution scenarios, such as PPA and Mirror. Accordingly, we conduct a further assessment of MI attacks under an increased resolution of , which is considered a more challenging task. The evaluation results for white-box and black-box attacks are demonstrated in Table 4.
The results imply the significant impact of GAN priors when attacking private images with higher resolution. All the methods that employ stronger GAN priors maintain low FID scores, including Mirror, C2FMI, and PPA, while other methods suffer from significant degradation in image reality. This phenomenon is more pronounced in the visualization results displayed in Fig 3. Despite the primary metric for evaluating MI attacks is Accuracy, the reality of reconstructed images should be ensured within a reasonable range for better image quality. Thus, it is imperative to explore more complex GAN priors with enhanced performance in future research, extending the MI field to more challenging and practical applications.
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror(white) | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI | |||||
| BREPMI | |||||
| Mirror (black) | |||||
| C2FMI | 47.317 | ||||
| LOKT |
4.4 Evaluation on Different Model Predictive Power
The predictive power of the target model is a crucial factor in determining the effectiveness of MI attacks. Previous work (Zhang et al., 2020a) has conducted preliminary experimental validations on simple networks (e.g. LeNet (LeCun et al., 1989)), demonstrating that the performance of the first GAN-based attack GMI is influenced by the predictive power of the target model. Therefore, we evaluate the state-of-the-art MI attacks on target models with varied predictive power to further validate the consistency of this characteristic in the recent attacks. The resolution of datasets is set to . The evaluation is stated in Fig. 4.
The comparison in Fig. 4 reveals most MI attacks maintain the trend that higher predictive power contributes to better attack performance, which is consistent with the aforementioned characteristic. Specifically, the earlier attack methods (e.g., GMI and KEDMI) presents more fluctuation on the trend across different predictive power, while the recent attack methods (e.g., PLGMI and PPA) show in more stable trend when predictive power increases. This indicates the predictive power of target models plays a crucial role in measuring the performance of MI attacks. Thus one can expect lower privacy leakage in the robust model by balancing the accuracy-privacy trade-off. This study of the predictive power of target models illustrates another useful aspect of our MIBench.
4.5 Evaluation on Defense Strategies
This analysis concentrates on the robustness of MI attacks when applied to target models with defense strategies. Notably, we select the first 100 classes subset from FaceScrub as the target dataset due to the time cost. The assessment results are listed in Fig.5. The configuration of defenses is set following the official parameters, as detailed in Table 8.
Overall, the TL (Ho et al., 2024) achieves the state-of-the-art decrease in Accuracy. However, advanced attack methods have overcome current defense strategies to some extent, such as PLG and PPA. Additionally, some older defense strategies (e.g. MID (Wang et al., 2021c)) are no longer effective against the latest attacks. From Fig.5, we observe that LS (Struppek et al., 2023) exhibits unexpectedly poor performance in metric while it was recently published in the top-tier conference. The potential reason might be the utilized target classifier with relatively high test accuracy, as validated in the Sec. C.4. Furthermore, we conduct further experiments with PPA as the attack method against ResNet-152 trained on high resolution (224×224) scenario, proving that this phenomenon can be extended into other defenses. The results are listed in the Table 5, demonstrating that all the defenses become invalid even with the recommended parameters. More in-depth evaluation on defenses is exhibited in Sec C.5.
Combining this phenomenon with the above experiment on predictive power, our empirical analysis indicates that the leaked information is strongly correlated to the model prediction accuracy and current defenses cannot effectively reduce the privacy information without sacrifice of model performance. Our findings emphasize that more reliable and stable defense strategies should be studied due to the fact that high model prediction accuracy is crucial for application of AI technology.
| Method | Hyperparameters | Test Acc | |||||
| NO Defense | - | 98.510 | 0.972 | 0.990 | 307.714 | 0.588 | 50.259 |
| MID | 96.760 | 1.000 | 1.000 | 273.687 | 0.517 | 49.239 | |
| 95.680 | 0.990 | 1.000 | 276.889 | 0.526 | 51.227 | ||
| BiDO | 98.030 | 0.986 | 0.996 | 306.827 | 0.554 | 50.943 | |
| 97.430 | 0.968 | 0.996 | 320.191 | 0.594 | 50.132 | ||
| TL | 97.620 | 0.994 | 1.000 | 282.394 | 0.513 | 53.023 | |
| 97.530 | 0.982 | 0.996 | 306.528 | 0.561 | 51.489 |
5 Conclusion
In this paper, we develop MIBench, a comprehensive, unified, and reliable benchmark, and provide an extensible and reproducible toolbox for researchers. To the best of our knowledge, this is the first benchmark and first open-source toolbox in the MI field. Our benchmark encompasses 16 of the state-of-the-art MI attack methods and defense strategies and more algorithms will be continually updated. Based on the implemented toolbox, we establish a consistent experimental environment and conducted extensive experimental analyses to facilitate fair comparison between different methods. In our experiments, we explore the impact of multiple settings, such as different image resolutions, model predictive power and defense. With in-depth analysis, we have identified new insights and proposed potential solutions to alleviate them.
Societal Impact and Ethical Considerations. A potential negative impact of our benchmark could be malicious users leveraging the implemented attack methods to reconstruct private data from public system. To alleviate this potential dilemma, a cautious approach for data users is to adopt robust and reliable defense strategies, as shown in the Sec.4.5 of our paper. Additionally, establishing access permissions and limiting the number of visits for each user is crucial to build responsible AI systems, thereby alleviating the potential contradictions with individual data subjects.
Appendix Table
Appendix A Benchmark Details
A.1 Details of Data Processing
This section introduces three types of dataset processing implemented by our toolbox, including data pre-processing, dataset splitting and dataset synthesis.
A.1.1 Data Pre-processing.
Data pre-processing aligns face images to ensure consistency across datasets. Users can customize the transformations to be used for data pre-processing. We also provide default processing for four datasets in high and low versions of the resolution. Low resolution versions include and , and high resolution versions include and . Here are the default pre-processing method for the four datasets.
FaceScrub (Ng & Winkler, 2014). In the low-resolution version, the original data is center cropped at and then scaled, and the high-resolution is scaled directly.
A.1.2 Dataset Splitting.
This step is used to model training, including classifiers and conditional generators. For labeled datasets like CelebA and FaceScrub, we provide a fixed partition to slice the training and test sets. For unlabeled datasets such as FFHQ and MetFaces, we provide a script for pseudo-labeling all images by computing the scores. The strategy of PLGMI (Yuan et al., 2023a) for the score calculation is contained in the examples of our toolbox.
A.1.3 Dataset Synthesis.
Some methods use synthetic datasets for model training. We provide a script to synthesis datasets by pre-trained GANs according to LOKT (Nguyen et al., 2023a).
A.2 Implementation of Attack Models.
We implemented and versions for all the custom attack models defined by the official code of implemented algorithms. For those algorithms that use StyleGAN2 (Karras et al., 2020b), we provide a wrapper for loading the official model and adapting it to our attack process.
A.3 Implementation of Classifier Training & Defense Methods
Our toolbox supports the training of a wide range of classifiers, including those used by the official code of most implemented algorithms, as well as those supported by TorchVision.
To ensure consistency across different defense strategies, we provide a unified framework for classifier training. Here we present the general idea of implemented methods.
1.
NO Defense. The classifiers are trained with the cross-entropy loss function.
2.
MID (Wang et al., 2021c). It applies a Gaussian perturbation to the features in front of the classification head to reduce the mutual information of the model inputs and outputs. According to the official code, the method can also be called VIB, and we adopt this name in our toolbox. The hyperparameter controls the factor of the Gaussian perturbation term.
3.
BiDO (Peng et al., 2022). It adds the regular term loss (COCO or HSIC) function so that the intermediate features decrease the mutual information with the input and increase the mutual information with the output. It has two hyperparameters, and . The former is the factor of the regular term loss between inputs and features, and the latter means that between features and model outputs.
4.
Label Smoothing (LS) (Struppek et al., 2023). It adds an label smoothing term with a negative smoothing factor into the cross-entropy loss.
5.
Transfer Learning (TL) (Ho et al., 2024). This method uses a pre-trained model for fine-tuning. The parameters of the previous layers are frozen and those of the later layers will be fine-tuned. In our implementation, the hyperparameter is defined as a ratio of the number of frozen layers.
In order to fairly compare the various defense methods, all defense methods are trained using the same pre-trained model in our experiments.
A.4 Details of Attack Process
The attack process follows a sequential workflow, containing latent vectors sampling, labeled latent vectors selection, optimization and final images selection. Here are the details of the workflow.
Latent Vectors Sampling.
This step generate random latent vectors and distributes them for each attack target. Most attack methods use a random distribution strategy. Mirror, PPA and BREP calculate the score of each latent vector corresponding to each label, and for each label the vectors with the highest scores are selected.
Labeled Latent Vectors Selection (Optional).
The previous step distributes latent vectors for each label, and this step optimizes the latent vectors by calculating the scores of the latent vectors corresponding to the labels and selecting the few vectors with the highest scores. Although currently there are no algorithms use this step, it can be added into the attack algorithms that use conditional generators, e.g. PLGMI (Yuan et al., 2023a) and LOKT (Nguyen et al., 2023a).
Optimization.
Optimization is the key step for the attack process, it accepts the initialized latent vectors and attack labels as input and outputs the optimized inverted images. We provide several kinds of optimization strategies of each attack method as follows.
- 1.
2.
Variance White-Box Optimization. It optimizes results from a Gaussian distribution of latent vectors corresponding to the target labels, and images are generated by random sampling from the optimized distribution. It is used by KEDMI (Chen et al., 2021b).
- 3.
- 4.
5.
BREP Optimization. A specific optimizer for BREPMI algorithm (Kahla et al., 2022). It uses a boundary repelling strategy for gradient simulation.
6.
Reinforcement Learning Optimization. Optimizing the latent vectors via reinforcement learning. Used by RLB attack method (Han et al., 2023).
Final Images Selection (Optional).
This step works by calculating the scores of each image and selecting the part of the image with the highest score as the result of the attack. It is used by PPA (Struppek et al., 2022).
| Method | Latent Optimizer | Identity Loss | Prior Loss | Image Augment |
| GMI | Momentum SGD | CE | Discriminator | ✗ |
| KEDMI | Adam | CE | Discriminator | ✗ |
| Mirror | Adam | CE | - | ✔ |
| PPA | Adam | Poincaré | - | ✔ |
| PLGMI | Adam | Max Margin | - | ✔ |
| LOMMA | Adam | Logit | Feature Distance | ✗ |
| IF-GMI | Adam | Poincaré | - | ✔ |
| LOKT | Adam | Max Margin | - | ✔ |
A.5 Details of Evaluation
We provide the following four evaluation metrics to evaluate the effectiveness of the attack.
1.
Classification Accuracy. The metric uses a given classifier to classify the inverted images and measures the top- and top- accuracy for target labels. The higher the reconstructed samples achieve attack accuracy on another classifier trained with the same dataset, the more private information in the dataset can be considered to be exposed (Zhang et al., 2020a).
2.
Feature Distance. The feature is defined as the output of the given classifier’s penultimate layer. We compute the shortest feature distance between inverted images and private samples for each class and calculate the average distance. Smaller feature distance means more similar features to the private image.
- 3.
- 4.
Appendix B Experimental Details
This section describes the setup and details of the experiments in this paper.
B.1 Experiment Settings
We conducted experiments at both low and high resolution scenarios. For the low resolution experiments, we employed classifiers with a resolution of as the target models, and an ResNet- with a resolution of served as the evaluation model. In the high resolution experiments, we used classifiers with a resolution of as the target models, with an Inception-v3 model having a resolution of as the evaluation model. Additionally, the computation resources utilized in our experiments including NVIDIA RTX 4090 and NVIDIA A100.
B.2 Datasets
The datasets used in our experiments are categorized into two types: public datasets and private datasets. The private datasets are used to train the target and evaluation models. Specifically, we selected identities with the most images from the CelebA dataset and all identities from the FaceScrub dataset as our private datasets.
The public datasets serve as a priori knowledge for the attacker to train the generator or to extract features of real faces. For the low-resolution experiments, we used FFHQ and the images from the CelebA dataset that are not included in the private dataset as our public datasets. For the high-resolution experiments, FFHQ and MetFaces were chosen as the public datasets. Note that MetFaces is an image dataset of 1336 human faces extracted from the Metropolitan Museum of Art Collection. It has a huge distribution shift with real human faces, which makes model inversion attack algorithms encounter great challenges.
The preprocessing of these datasets is described in A.1.1.
B.3 Classifiers
For the attack experiments, we trained multiple classifiers as target models and evaluation models, as detailed in Table 7. For the defense experiments, we trained the ResNet- model on the FaceScrub dataset using various defense methods, as outlined in Table 8.
| Dataset | Model | Resolution | Test Acc |
| CelebA | VGG-16 | 0.8826 | |
| ResNet-152 | 0.9371 | ||
| ResNet-50 | 0.9588 | ||
| ResNet-152 | 0.9003 | ||
| Inception-v3 | 0.9216 | ||
| FaceScrub | VGG-16 | 0.8785 | |
| ResNet-152 | 0.9825 | ||
| ResNet-50 | 0.9938 | ||
| ResNet-152 | 0.9225 | ||
| ResNeSt-101 | 0.9329 | ||
| Inception-v3 | 0.9445 |
| Defense Method | Hyperparameters | Test Acc |
| No Defense | - | |
| MID | ||
| BiDO | ||
| LS | ||
| TL |
B.4 Evaluation.
The definitions of the evaluation metrics are detailed in Section A.5. Here, we present the specific details of the metrics used in the experiments.
For the Classification Accuracy and Feature Distance metrics, we evaluate the attack results using another classifier pre-trained on the same dataset as the target model: ResNet-50 for low-resolution settings and Inception-v3 for high-resolution settings, denoted as Acc and . Additionally, an Inception-v1 model pre-trained on a large face dataset, VGGFace2, is used to calculate the feature distance, measuring the realism of the inverted images, denoted as .
For FID, Precision-Recall, and Density-Coverage scores, we follow the experimental setup of existing papers. We use Inception-v3, pre-trained on ImageNet, to extract the features of images and participate in the score calculation.
Appendix C More Experimental Results
C.1 Sample Diversity
Following the settings of Section 4.2 and 4.3, we computed the Precision-Recall (Kynkäänniemi et al., 2019) and Density-Coverage (Kynkäänniemi et al., 2019) to evaluate the intra-class diversity for each attack method. The results are presented in Table 9, 10, 11 and 12. It tends to be that the attacks with stronger GAN priors get higher scores, such as Mirror, C2FMI and PPA.
| Method | ||||
| GMI | ||||
| KEDMI | ||||
| Mirror(white) | ||||
| PPA | ||||
| PLGMI | ||||
| LOMMA+GMI | ||||
| LOMMA+KEDMI | ||||
| IF-GMI |
| Method | ||||
| BREP | ||||
| Mirror(black) | ||||
| C2F | ||||
| LOKT |
| Method | ||||
| GMI | ||||
| KEDMI | ||||
| Mirror(white) | ||||
| PPA | ||||
| PLGMI | ||||
| LOMMA+GMI | ||||
| LOMMA+KEDMI | ||||
| IF-GMI |
| Method | ||||
| BREP | ||||
| Mirror(black) | ||||
| C2F | ||||
| LOKT |
C.2 Evaluation on Different Target Classifiers
In addition to attacking IR-152 and ResNet-152 in Section 4.2 and 4.3, we extend our experiments on more different target classifiers. We evaluate attacks on VGG- (Simonyan & Zisserman, 2014) and ResNeSt- (Zhang et al., 2022a) in low and high resolution settings, respectively. The results are presented in Table 13 and 14.
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror(white) | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI |
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror(white) | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI |
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI |
Besides the aforementioned CNN-based classifiers, we further analyze MI attacks on ViT-based models. Table 15 presents further experiments on the MaxViT (Tu et al., 2022) under the high resolution (224×224) scenario. Other experimental settings are continuous with the main paper, with FFHQ as the public dataset and FaceScrub as the private dataset. The test accuracy of the target MaxViT is 94.61%.
C.3 Evaluation on More Combination of Datasets
Evaluations in the Section 4 are conducted under the same dataset combination of FFHQ as the public dataset and FaceScrub as the private dataset. Therefore, we design more combination of datasets in this part to further assess the transferability of different attacks. The results are listed in Table 16 and 17. The visual results are shown in Figure 6 and 7.
Except for the typical face classification task, we have conducted more experiments on the dog breed classification task under the high resolution (224×224) scenario, which includes two non-facial datasets, Stanford Dogs (Dataset, 2011) and Animal Faces-HQ Dog (AFHQ) (Choi et al., 2020). The public dataset is AFHQ while the private dataset is Stanford Dogs. The target model is ResNet-152 of 77.45% test accuracy, which follows the same setting in PPA. Table 18 lists the evaluation results.
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror(white) | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI |
| Method | |||||
| GMI | |||||
| KEDMI | |||||
| Mirror(white) | |||||
| PPA | |||||
| PLGMI | |||||
| LOMMA+GMI | |||||
| LOMMA+KEDMI | |||||
| IF-GMI |
| Method | ||||
| GMI | ||||
| KEDMI | ||||
| Mirror | ||||
| PPA | ||||
| PLGMI | ||||
| LOMMA+GMI | ||||
| LOMMA+KEDMI | ||||
| IF-GMI |
C.4 Evaluation on LS Defense Method
To evaluate the LS defense method, we follow the settings in the official paper (Struppek et al., 2023) and evaluate it in more settings, shown in Table 19. To better evaluate the effectiveness of LS defense algorithms by making the undefended classifiers slightly less accurate than the defense-imposed classifiers through an early-stop strategy when training.
In setting A, the public and private datasets are different part of CelebA dataset in low resolution scenario with the GMI as the attack method. In setting B, the public dataset is FFHQ and the private dataset is FaceScrub with the PPA as the attack method. In these two settings, the predictive power is much lower than that describes in Section 4.5. In this case, the LS defense is very effective, making the success rate of the attack drop dramatically.
| Setting | Defense | Test Acc | ||||
| A | - | |||||
| LS | ||||||
| B | - | |||||
| LS |
C.5 Validation for Defense Methods on Models with Low Accuracy
The results for models with low accuracy are listed in Table 20. Obviously, the defense is valid when the target model has relatively low prediction accuracy.
| Method | Hyperparameters | Test Acc | |||||
| NO Defense | - | 92.170 | 0.686 | 0.914 | 262.471 | 0.767 | 68.454 |
| MID | 88.240 | 0.568 | 0.820 | 246.021 | 0.757 | 69.663 | |
| BiDO | 88.620 | 0.582 | 0.874 | 275.453 | 0.793 | 68.248 | |
| TL | 89.160 | 0.316 | 0.616 | 279.439 | 0.897 | 63.241 |
C.6 Evaluation on Different Loss Functions
In recent years, various attack algorithms have attempted to mitigate the effects of gradient vanishing by employing different loss functions. In this part, we investigate the impact of identity loss functions on the success rate of model inversion attacks. Specifically, we adopt PPA with FFHQ prior to attack a ResNet-152 classifier pre-trained on FaceScrub. The comparison of the results is presented in Table 21. Our findings indicate that the Poincare loss function yields the highest performance without model augmentation, whereas the Logit loss function achieves the best results with model augmentation.
| Loss Function | |||||
| CE | |||||
| Poincaré | |||||
| Max Margin | |||||
| Logit | |||||
| CE+ | |||||
| Poincaré+ | |||||
| Max Margin+ | |||||
| Logit+ |
C.7 More evaluation on VMI, RLBMI and PPA
Considering the high computational overhead of RLBMI and VMI, we only experiment at a low resolution settings for classes. The public and private datasets are different part of CelebA dataset. The results are shown in Table 22.
| Method | ||||
| VMI | ||||
| RLBMI |
We also explored the effect of PPA for different number of latent vectors for optimization and the number of iterations. It is presented in Table 23. Note that PPA select top- optimized latent vectors as attack results.
| Number of latents | Iterations | |||||
| 20 | 50 | |||||
| 50 | 50 | |||||
| 20 | 600 |
Appendix D Limitations and Future Plans
Our benchmark mainly focuses on GAN-based MI attacks and MI defenses applied in the classifier training stage. We will extend our benchmark to wider range of MI methods, including learning-based MI attacks and MI defenses applied to the classifier output. Furthermore, the concept of MI also pervades across modalities (e.g. text (Parikh et al., 2022; Zhang et al., 2023; Carlini et al., 2019; 2021; Yu et al., 2023; Nasr et al., 2023) and graph learning (Zhang et al., 2021; 2022b; Zhou et al., 2023; Wu et al., 2022; He et al., 2021)) beyond computer vision domain, where our benchmark concentrates in current stage. Expansion to new modalities is a promising direction for our benchmark to further explore the privacy threats in other fields, leading to more generalization in the AI security. In addition to developing new algorithms, it is also essential to conduct further research on the MI attacks and defenses to make in-depth analysis about their characteristics and bring valuable new insights.

### v3

@@
-MIBench: A Comprehensive Benchmark for Model Inversion Attack and Defense
+MIBench: A Comprehensive Framework for Benchmarking Model Inversion Attack and Defense
 Abstract
-Model Inversion (MI) attacks aim at leveraging the output information of target models to reconstruct privacy-sensitive training data, raising widespread concerns on privacy threats of Deep Neural Networks (DNNs).
-Unfortunately, in tandem with the rapid evolution of MI attacks, the lack of a comprehensive, aligned, and reliable benchmark has emerged as a formidable challenge.
-This deficiency leads to inadequate comparisons between different attack methods and inconsistent experimental setups.
-In this paper, we introduce the first practical benchmark for model inversion attacks and defenses to address this critical gap, which is named MIBench.
-This benchmark serves as an extensible and reproducible modular-based toolbox and currently integrates a total of 16 state-of-the-art attack and defense methods.
-Moreover, we furnish a suite of assessment tools encompassing 9 commonly used evaluation protocols to facilitate standardized and fair evaluation and analysis.
-Capitalizing on this foundation, we conduct extensive experiments from multiple perspectives to holistically compare and analyze the performance of various methods across different scenarios, which overcomes the misalignment issues and discrepancy prevalent in previous works.
-Based on the collected attack methods and defense strategies, we analyze the impact of target resolution, defense robustness, model predictive power, model architectures, transferability and loss function.
-Our hope is that this MIBench could provide a unified, practical and extensible toolbox and is widely utilized by researchers in the field to rigorously test and compare their novel methods, ensuring equitable evaluations and thereby propelling further advancements in the future development.
+Model Inversion (MI) attacks aim at leveraging the output information of target models to reconstruct privacy-sensitive training data, raising critical concerns regarding the privacy vulnerabilities of Deep Neural Networks (DNNs).
+Unfortunately, in tandem with the rapid evolution of MI attacks, the absence of a comprehensive benchmark with standardized metrics and reproducible implementations has emerged as a formidable challenge.
+This deficiency has hindered objective comparison of methodological advancements and reliable assessment of defense efficacy.
+To address this critical gap, we build the first practical benchmark named MIBench for systematic evaluation of model inversion attacks and defenses.
+This benchmark bases on an extensible and reproducible modular-based toolbox which currently integrates a total of 19 state-of-the-art attack and defense methods and encompasses 9 standardized evaluation protocols.
+Capitalizing on this foundation, we conduct extensive evaluation from multiple perspectives to holistically compare and analyze various methods across different scenarios, such as the impact of target resolution, model predictive power, defense performance and adversarial robustness.
 1 Introduction
-In recent years, Model Inversion (MI) attacks have raised alarms over the potential privacy breaches of sensitive personal information, including the leakage of privacy images in face recognition models (He et al., 2016), sensitive health details in medical data (Wang et al., 2022), financial information such as transaction records and account balances (Ozbayoglu et al., 2020), and personal preferences and social connections in social media data (Feng et al., 2022).
-In the MI attacks, an attacker aims to infer private training data from the output information of the target model.
-Fredrikson et al.
-(2014) proposed the first MI attack against linear regression models to reconstruct sensitive features of genomic data.
-Subsequent studies (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) have extended MI attacks to more Machine Learning (ML) models.
-Zhang et al.
-(2020b) first introduces the GANs as stronger image priors, paving the way for more applications in GAN-based methods (Chen et al., 2021a; Yuan et al., 2022; 2023b; Wang et al., 2021b).
-However, some critical challenges are posed owing to the lack of a comprehensive, fair, and reliable benchmark.
-Specifically, the evaluation of new methods is often confined to comparisons with a narrow selection of prior works, limiting the scope and depth of analysis.
-For instance, some methods (Nguyen et al., 2023b; Yuan et al., 2023a) exhibit superior performance for lower-resolution images while other methods (Struppek et al., 2022; Qiu et al., 2024) perform better at higher resolutions.
-However, these studies only evaluate under their predominant resolutions, and thus do not provide a unified and comprehensive comparison.
-Additionally, the absence of unified experimental protocols results in a fragmented landscape where there is less validity and fairness in the comparative studies.
-For example, the original GMI (Zhang et al., 2020a) achieves the attack accuracy of 28%, 44%, 46% when attacking three different classifiers trained on the CelebFaces Attributes(CelebA) (Liu et al., 2015), while merely maintaining lower attack accuracy of 21%, 32%, 31% and 21%, 31%, 29% when compared in the KEDMI (Chen et al., 2021b) and PLGMI (Yuan et al., 2023a) under the same experimental setup.
-Compounding these issues, discrepancies in the evaluation metrics further obscure the reliability of reported conclusions, potentially steering the field towards biased or misleading insights.
+Model Inversion (MI) attacks aim to infer private training data from the output information of target models.
+In recent years, the proliferation of MI attacks have raised alarms over the potential privacy breaches of sensitive personal information, including the leakage of privacy images in face recognition models [17], sensitive health details in medical data [51], financial information such as transaction records and account balances [36], and personal preferences and social connections in social media data [10].
+Since [64] first introduced the Generative Adversarial Networks (GANs) as stronger image priors and established the first GAN-based MI framework, a series of subsequent studies [2, 61, 50, 39] have driven substantial progress in this emerging field during the past few years.
+Despite the remarkable development in the MI field, some critical challenges are posed due to the absence of a comprehensive and reliable benchmark.
+1) Limited Reproducibility: While most MI methods provide open-source implementations, significant reproducibility challenges persist.
+First, heterogeneous runtime environments [43, 60, 49] impede simultaneous replication across different studies.
+Second, the unavailability of pre-trained models in some works [63, 3, 49] hinders comprehensive comparative experimentation.
+2) Inadequate and Unfair comparison on different methods: The evaluation of new methods is often confined to comparisons with a narrow selection of prior works, limiting the scope and depth of analysis.
+Some methods [35, 60] exhibit superior performance for lower-resolution images while other methods [43, 39] perform better at higher resolutions.
+However, these studies only perform evaluation under their predominant resolutions, and thus do not provide a unified and comprehensive comparison.
 These shortcomings impede both the accurate measurement of advancements in the MI field and the systematic exploration of its theoretical underpinnings, underscoring the urgent need for a harmonized framework to facilitate robust and transparent research practices.
-To alleviate these problems, we establish the first benchmark in MI field, named MIBench.
-For building an extensible modular-based toolbox, we disassemble the pipeline of MI attacks and defenses into four main modules, each designated for data preprocessing, attack methods, defense strategies and evaluation, hence enhancing the extensibility of this unified framework.
-The proposed MIBench has encompassed a total of 16 distinct attack and defense methods, coupled with 9 prevalent evaluation protocols to adequately measure the comprehensive performance of individual MI methods.
-Furthermore, we conduct extensive experiments from diverse perspectives to achieve a thorough appraisal of the competence of existing MI methods, while simultaneously venturing into undiscovered insights to inspire potential avenues for future research.
-We expect that this reproducible benchmark will facilitate the further development of MI field and bring more innovative explorations in the subsequent study.
+To alleviate these problems, we establish the first comprehensive benchmark named MIBench for MI attacks and defenses.
+Our proposed MIBench is built upon an extensible modular-based toolbox, which has encompassed a total of 19 distinct attack and defense methods and coupled with 9 prevalent evaluation protocols to adequately measure the comprehensive performance of individual MI methods.
+Furthermore, we conduct extensive evaluation from multiple perspectives to achieve a thorough appraisal for existing MI methods, while simultaneously venturing into undiscovered insights to inspire potential avenues for future research.
+We expect that this reproducible and reliable benchmark will facilitate the further development of MI field and bring more innovative explorations in the subsequent study.
 Our main contributions are as follows:
-We build the first comprehensive benchmark in MI field, which serves as an extensible and reproducible modular-based toolbox for researchers.
-We expect that this will foster the development of more powerful MI attacks by making it easier to evaluate their effectiveness across multiple distinct dimensions.
-We implement 16 state-of-the-art attack methods and defense strategies and 9 evaluation protocols currently in our benchmark.
-We hope that the benchmark can further identify the most successful ideas in defending against rapid development of potential MI attacks.
-We conduct extensive experiments to thoroughly assess different MI methods under multiple settings and study the effects of different factors to offer new insights on the MI field.
-In particular, we validate that stronger model predictive power correlates with an increased likelihood of privacy leakage.
-Moreover, our analysis reveals that certain defense algorithms also fail when the target model achieves high prediction accuracy.
+We build the first comprehensive benchmark in MI field, featuring a reliable and reproducible modular-based toolbox and comprehensive evaluation on multiple scenarios.
+We currently implement a total of 19 state-of-the-art attack methods and defense strategies and 9 standarized evaluation protocols in our toolbox.
+We conduct extensive experiments to thoroughly assess different MI attacks and defenses under multiple settings and investigate the effects of different factors to offer new insights on the MI field.
 2 Related Work
 Model Inversion Attacks.
-In the MI attacks, the malicious adversary aims to reconstruct privacy-sensitive data by leveraging the output prediction confidence of the target classifier and other auxiliary priors.
-Normally, the attacker requires a public dataset that shares structural similarities with the private dataset but without intersecting classes to pre-train the generator.
+In MI attacks, the malicious adversary aims to reconstruct privacy-sensitive data by leveraging the output prediction confidence of target classifiers and other necessary auxiliary priors.
+Normally, the attacker requires a public dataset that shares structural similarities but has no overlapping labels with the private dataset, which is utilized to pre-train a generator before launching an attack.
 For example, an open-source face dataset serves as essential public data when targeting a face recognition model.
-For a typical GAN-based MI attack, attackers attempt to recover private images from the GAN’s latent vectors initialized by Gaussian distribution, given the target image classifier parameterized with weights and the trained generator .
-The attack process can be formulated as follows:
-where is the target class, typically denotes the classification loss, is a hyperparameter, and is the prior knowledge regularization (e.g., the discriminator’s classification loss) used to improve the reality of .
-By minimizing the above equation, the adversary updates the latent vectors into the optimal results and generate final images through .
-Fredrikson et al.
-(2014) first introduce the concept of MI attacks in the context of genomic privacy.
-They find that maximizing the posterior probabilities of a linear regression model can reconstruct the original genomic markers.
-Ensuing works (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) manage to design MI attacks for more kinds of models and private data, but are still limited to attacking simple networks and grayscale images.
-To enhance the reconstruction performance on complex RGB images, GMI (Zhang et al., 2020a) first propose to incorporate the rich prior knowledge (Fang et al., 2023; Gu et al., 2020; Fang et al., 2024) within the pre-trained Generative Adversarial Networks (GANs) (Goodfellow et al., 2014).
+In typical MI attacks based on Generative Adversarial Networks (GANs), attackers attempt to recover private images from the GAN’s latent vectors initialized by Gaussian distribution, given the target image classifier parameterized with weights and the trained generator .
+The attack process is formulated as follows:
+where is the target class, typically denotes the classification loss, is a hyperparameter, and is the prior knowledge regularization (e.g., the discriminator’s classification loss) used to improve the image fidelity of .
+Following the above optimization, the adversary updates the latent vectors into the optimal results and generate final images through .
+Based on the accessibility of target models, MI attacks can be further split into white-box attacks, black-box attacks, and label-only attacks.
+In white-box settings, the full parameters of target models are available to attackers.
+GMI [63] first proposes to incorporate the rich prior knowledge [8, 15, 9] within the pre-trained GANs [13].
 Specifically, GMI starts by generating a series of preliminary fake images, and then iteratively optimizes the input latent vectors that are used for generation.
-Based on GMI, KEDMI (Chen et al., 2021b) refine the discriminator by introducing target labels to recover the distribution of the input latent vectors.
-VMI (Wang et al., 2021a) utilize variational inference to model MI attacks and adopts KL-divergence as the regularization to better estimate the target distribution.
-PPA (Struppek et al., 2022) introduce a series of techniques such as initial selection, post-selection, and data argumentation to enhance MI attacks and manages to recover high-resolution images by the pre-trained StyleGAN2 (Karras et al., 2019).
-LOMMA (Nguyen et al., 2023b) integrate model augmentation and model distillation into MI attacks to tackle the problem of over-fitting.
-PLGMI (Yuan et al., 2023a) leverage a top- selection technique to generate pseudo labels to further guide the training process of GAN.
-Besides, based on whether the parameters and structures of the victim model are fully accessible to the attackers, MI attacks can be further split into white-box attacks and black-box attacks.
-Note that in black-box settings, the gradients can no longer be computed by the back-propagation process.
-Thus, Yuan et al.
-(2022) address this problem by sampling numerous latent vectors from random noise, selecting the ones that can generate correct labels, and updating the latent vectors solely on discriminator loss.
-Nguyen et al.
-(2023a) propose to conduct MI attacks on various surrogate models instead of the unfamiliar victim model, transforming the black-box settings into white-box settings.
+Based on GMI, KEDMI [3] refines the discriminator by introducing target labels to recover the distribution of the input latent vectors.
+VMI [49] utilizes variational inference to model MI attacks and adopts KL-divergence as the regularization to better estimate the target distribution.
+PPA [43] introduces a series of techniques such as initial selection, post-selection, and data argumentation to enhance MI attacks and manages to recover high-resolution images by the pre-trained StyleGAN2 [22].
+LOMMA [35] integrates model augmentation and model distillation into MI attacks to tackle the problem of over-fitting.
+PLGMI [60] leverages a top- selection technique to generate pseudo labels to further guide the training process of GAN.
+The most recent work IFGMI[39] delves into the internal architecture of GANs and achieve higher-quality image inversion through optimization of the intermediate-layer feature vectors.
+In black-box settings, the attackers only have access to the output prediction confidence vectors (also named soft-label).
+Thus, gradients of the target model can no longer be computed by the back-propagation process.
+Mirror [1] addresses this problem by employing a genetic algorithm as an alternative to gradient descent for optimization.
+Following [1], C2FMI [59] designs an inverse net to map soft labels into the latent space of pre-trained GANs, which serves as coarse optimization before launching the genetic algorithm.
+RLBMI [16] presents a novel paradigm by incorporating reinforcement learning as the optimization mechanism.
+The label-only settings step further on the black-box settings, relaxing the accessibility to merely one-hot vectors containing the final predicted class (also named hard-label).
+BREPMI [21] estimates the optimization direction via the boundary repulsion technique.
+LOKT[34] proposes to conduct MI attacks on various surrogate models instead of the unfamiliar victim model, transforming the label-only settings into white-box settings.
 Model Inversion Defenses.
-To defend MI attacks, most existing methods can be categorized into two types: model output processing (Yang et al., 2020; Wen et al., 2021; Ye et al., 2022) and robust model training (Gong et al., 2023; Titcombe et al., 2021; Li et al., 2022; Wang et al., 2021c; Peng et al., 2022; Struppek et al., 2023).
-Model output processing refers to reducing the private information carried in the victim model’s output to promote privacy.
-Yang et al.
-(2020) propose to train an autoencoder to purify the output vector by decreasing its degree of dispersion.
-Wen et al.
-(2021) apply adversarial noises to the model output and confuse the attackers.
-Ye et al.
-(2022) leverage a differential privacy mechanism to divide the output vector into multiple sub-ranges.
-Robust model training refers to that incorporating the defense strategies during the training process.
-MID Wang et al.
-(2021c) penalizes the mutual information between model inputs and outputs in the training loss, thus reducing the redundant information carried in the model output that may be abused by the attackers.
+To defend MI attacks, most existing methods can be categorized into two types according to their functional stage: train-time defense [12, 47, 29, 52, 37, 44] and inference-time defense [56, 53, 58].
+Train-time defense refers to incorporating the defense strategies during the training process.
+MID [52] proposes penalizing the mutual information between model inputs and outputs in the training loss to reduce the redundant information carried in the model output that may be abused by the attackers.
 However, simply decreasing the dependency between the inputs and outputs also results in model performance degradation.
-To alleviate this issue and strike a better balance between model utility and user privacy, Bilateral Dependency Optimization (BiDO) (Peng et al., 2022) minimizes the dependency between the inputs and outputs while maximizing the dependency between the latent representations and outputs.
-Gong et al.
-(2023) propose to leverage GAN to generate fake public samples to mislead the attackers.
-Titcombe et al.
-(2021) defend MI attacks by adding Laplacian noise to intermediate representations.
-LS (Struppek et al., 2023) finds that label smoothing with negative factors can help privacy preservation.
-TL (Ho et al., 2024) leverages transfer learning to limit the number of layers encoding sensitive information and thus improves the robustness to MI attacks.
+To strike a better balance between model utility and user privacy, BiDO [37] minimizes the dependency between the inputs and outputs while maximizing the dependency between the latent representations and outputs.
+[12] propose to leverage GAN to generate fake public samples to mislead the attackers.
+[47] defend MI attacks by adding Laplacian noise to intermediate representations.
+LS [44] finds that label smoothing with negative factors can help privacy preservation.
+TL [20] leverages transfer learning to limit the number of layers encoding sensitive information and thus improves the robustness to MI attacks.
+The most recent work [25] points out that the skip connection structure of DNNs enhances MI attacks and first proposes defenses from the perspective of robust model architecture.
+Inference-time defense refers to reducing the private information carried in the victim model’s output to promote privacy without modifying the training process.
+This type of defense often functions as an additional plugin.
+Purifier [57] proposed to train an autoencoder to purify the output vector by increasing its degree of dispersion.
+SSD [65] proposed to minimize the conditional mutual information under the constraint of bounded distortion.
+| Attack Method | Accessibility | Reference | GAN Prior | Official Resolution |
+| GMI [63] | White-box | CVPR-2020 | WGAN | |
+| KEDMI [3] | White-box | ICCV-2021 | Inversion-specific GAN† | |
+| VMI [49] | White-box | NeurIPS-2021 | StyleGAN2 | |
+| PPA [43] | White-box | ICML-2022 | StyleGAN2 | |
+| PLGMI [60] | White-box | AAAI-2023 | Conditional GAN | |
+| LOMMA [35] | White-box | CVPR-2023 | ||
+| IF-GMI [39] | White-box | ECCV-2024 | StyleGAN2 | |
+| Mirror∗ [1] | White-box/Black-box | NDSS-2022 | StyleGAN | |
+| C2FMI [59] | Black-box | TDSC-2023 | StyleGAN2 | |
+| RLBMI [16] | Black-box | CVPR-2023 | WGAN | |
+| BREPMI [21] | Label-only | CVPR-2022 | WGAN | |
+| LOKT [34] | Label-only | NeurIPS-2023 | ACGAN |
+| Defense Strategy | Functional Stage | Reference | Core Technique | Description |
+| MID [52] | Train-time | AAAI-2021 | Regularization | Utilize mutual information regularization to limit leaked information about the model input in the model output |
+| BiDO [37] | Train-time | KDD-2022 | Regularization | Minimize dependency between latent vectors and the model input while maximizing dependency between latent vectors and the outputs |
+| LS [44] | Train-time | ICLR-2024 | Label Smoothing | Adjusting the label smoothing with negative factors contributes to increasing privacy protection |
+| TL [20] | Train-time | CVPR-2024 | Transfer Learning | Utilize transfer learning to limit the number of layers encoding privacy-sensitive information for robustness to MI attacks |
+| RoLSS/SSF/TTS∗ [25] | Train-time | ECCV-2024 | Skip Connection | Properly remove some skip connection structures in the target classifier during the training stage |
+| Purifier [57] | Inference-time | AAAI-2023 | Conditional VAE | Reform confidences and swap labels to make members behave like non-members |
+| SSD [65] | Inference-time | ICLR-2025 | Conditional Mutual Information | Minimize conditional mutual information to make predictions less dependent on inputs and more dependent on ground truths |
+∗The three defenses proposed by [25] have minor distinction between each other, therefore we select the TTS as the
+representative method in our evaluation.
 3 Our Benchmark
 3.1 Dataset
-Considering existing MI attacks primarily focus on reconstructing private facial data from image classifiers, we select 4 widely recognized face datasets as the basic datasets for our benchmark, which include Flickr-Faces-HQ (FFHQ) (Karras et al., 2019), MetFaces (Karras et al., 2020a), FaceScrub (Ng & Winkler, 2014), and CelebFaces Attributes (CelebA) (Liu et al., 2015).
+Considering existing MI attacks primarily focus on reconstructing private facial data from image classifiers, we select 4 widely recognized face datasets as the basic datasets, including Flickr-Faces-HQ (FFHQ) [22], MetFaces [23], FaceScrub [33], and CelebFaces Attributes (CelebA) [30].
 Generally, FFHQ and MetFaces are employed as public datasets for pre-training auxiliary priors, whereas FaceScrub and CelebA serve as the target private datasets to attack.
-Our benchmark facilitates researchers to freely combine public datasets with private ones, thereby enabling customized experimental setups.
-Extensive evaluation on more non-facial datasets are presented in Sec.C.3
+Our benchmark is modularized to facilitate researchers to freely combine public datasets with private ones, enabling customized experimental setups.
+Extensive evaluation on more non-facial datasets (Stanford Dogs [5] and AFHQ-Dog [4]) are presented in Supp.C.3.
 Notably, the target resolutions across different MI attacks are not uniform.
@@
 Therefore, our benchmark offers 2 versions of low-resolution and higher-resolution for the aforementioned 4 datasets and prepares multiple transformation tools for processing images, freeing researchers from the laborious tasks of data preprocessing.
-More details regarding the datasets can be found in the Sec.
-B.2 of the Appendix.
+More details regarding the datasets can be found in the Supp.B.2.
 3.2 Implemented Methods
-Our benchmark includes a total of 16 methods, comprising 11 attack methods and 4 defense strategies.
+Our benchmark includes a total of 19 methods, comprising 12 attack methods and 7 defense strategies.
 With a focus on Generative Adversarial Network (GAN)-based MI attacks, we selectively reproduce methods from recent years that have been published in top-tier conferences or journals in the computer vision or machine learning domains.
 This criterion ensures the reliability and validity of the implemented methods.
-Considering the main targets in our benchmark are image classifiers for RGB images, the learning-based MI attacks (Fredrikson et al., 2015; Song et al., 2017; Yang et al., 2019) are not incorporated currently.
-More detailed information about the implemented methods is stated in Sec.
-A.2 and Sec.
-A.3.
+Considering the main targets in our benchmark are image classifiers for RGB images, the learning-based MI attacks [11, 42, 55] are not incorporated currently.
+More detailed information about the implemented methods is stated in Supp.A.2 and Supp.A.3.
 Attacks.
@@
 Our benchmark includes 8 white-box attack methods and 4 black-box attack methods, as summarized in Table 1.
-| Attack Method | Accessibility | Reference | GAN Prior | Official Resolution |
-| GMI (Zhang et al., 2020a) | White-box | CVPR-2020 | WGAN-GP | |
-| KEDMI (Chen et al., 2021b) | White-box | ICCV-2021 | Inversion-specific GAN† | |
-| VMI (Wang et al., 2021a) | White-box | NeurIPS-2021 | StyleGAN2 | |
-| Mirror∗ (An et al., 2022) | White-box/Black-box | NDSS-2022 | StyleGAN | |
-| PPA (Struppek et al., 2022) | White-box | ICML-2022 | StyleGAN2 | |
-| PLGMI (Yuan et al., 2023a) | White-box | AAAI-2023 | Conditional GAN | |
-| LOMMA (Nguyen et al., 2023b) | White-box | CVPR-2023 | ||
-| IF-GMI (Qiu et al., 2024) | White-box | ECCV-2024 | StyleGAN2 | |
-| BREPMI (Kahla et al., 2022) | Black-box | CVPR-2022 | WGAN-GP | |
-| C2FMI (Ye et al., 2023) | Black-box | TDSC-2023 | StyleGAN2 | |
-| RLBMI (Han et al., 2023) | Black-box | CVPR-2023 | WGAN-GP | |
-| LOKT (Nguyen et al., 2023a) | Black-box | NeurIPS-2023 | ACGAN |
 Defenses.
 To effectively defend against MI attacks, the defender typically employs defense strategies during the training process of victim classifiers.
-Our benchmark includes 4 typical defense strategies and the details are presented in Table 2.
-| Defense Strategy | Reference | Core Technique | Description |
-| MID (Wang et al., 2021c) | AAAI-2021 | Regularization | Utilize mutual information regularization to limit leaked information about the model input in the model output |
-| BiDO (Peng et al., 2022) | KDD-2022 | Regularization | Minimize dependency between latent vectors and the model input while maximizing dependency between latent vectors and the outputs |
-| LS (Struppek et al., 2023) | ICLR-2024 | Label Smoothing | Adjusting the label smoothing with negative factors contributes to increasing privacy protection |
-| TL (Ho et al., 2024) | CVPR-2024 | Transfer Learning | Utilize transfer learning to limit the number of layers encoding privacy-sensitive information for robustness to MI attacks |
+Our benchmark has implemented 4 typical defense strategies and corresponding details are presented in Table 2.
 3.3 Toolbox
-We implement an extensible and reproducible modular-based toolbox for our benchmark, as shown in Fig 1.
-The framework can be divided into four main modules, including Data Preprocess Module, Attack Module, Defense Module and Evaluation Module.
+We implement an extensible and reproducible modular-based toolbox for our benchmark, as shown in Fig.
+1.
+The framework can be roughly divided into four primary modules, including Data Preprocess Module, Attack Module, Defense Module and Evaluation Module.
 Data Preprocess Module.
@@
 Consequently, we furnish this module with three fundamental functionalities: dataset preprocessing, target model training, and parameter management.
-For dataset preprocessing, we build a unified pipeline for each dataset, which automatically carries out a series of operations such as spliting dataset and image transformations (e.g.
+For dataset preprocessing, we build a unified pipeline for each dataset, which automatically carries out a series of operations such as splitting dataset and image transformations (e.g.
 center crop and resize) based on the split file and chosen resolution from users.
@@
 We abstract the general procedures in classifier training into a base trainer class to facilitate users in extending and customizing their own classifiers.
-For parameter management, we encapsulate parameters used in different processes into specific configuration classes, such as TrainConfig designated for the training process and AttackConfig for the attack process, thus ensuring organized and efficient parameter handling across various operations in the workflow.
+For parameter management, we encapsulate parameters used in different processes into specific configuration classes, such as TrainConfig, designated for the training process, and AttackConfig for the attack process, thus ensuring organized and efficient parameter handling in the workflow.
 Attack Module.
@@
 The second stage is the core inversion attack, which we split into three parts: latent vectors initialization, iterative optimization, and an optional post-processing step.
-After completion of the attack, the module preserves essential data such as the final optimized images and latent vectors, facilitating subsequent evaluation and analysis.
+After the completion of the attack, the module preserves essential data such as the final optimized images and latent vectors, facilitating subsequent evaluation and analysis.
 Defense Module.
@@
 Evaluation Module.
-Our benchmark concentrates on the evaluation at distribution level instead of sample level, assessing the overall performance of the whole reconstructed dataset.
-Therefore, we provide a total of 9 widely recognized distribution level evaluation metrics for users, which can be categorized into four types according to the evaluated content: Accuracy, Feature Distance, FID (Heusel et al., 2017) and Sample Diversity.
-For Accuracy, this metric measures how well the reconstructed samples resemble the target class, consisting of and .
-For Feature Distance, it computes the average shortest distance from features between each reconstructed sample and private data to measure the similarity between the feature space of reconstructed samples and private dataset.
-For FID, the lower FID score shows higher realism and overall diversity (Wang et al., 2021a).
-For Sample Diversity, the higher values indicate greater intra-class diversity.
-Moreover, we provide convenient tools for analysis, including Standard Deviation Calculation and Visualization.
-4 Experiment
-4.1 Experimental Setups
-To ensure fair and uniform comparison and evaluation, we select the FFHQ (Karras et al., 2019) as the public dataset and FaceScrub (Ng & Winkler, 2014) as the private dataset for all the experiments in the Experiment section.
-The target models are fixed to the IR-152 (He et al., 2016) for low-resolution scenario and ResNet-152 (He et al., 2016) for high-resolution scenario, both trained on the FaceScrub.
-For each attack method, the number of images reconstructed per class is set to 5 due to considerations of time and computation cost.
-More detailed experimental settings are listed in Sec.B in the Appendix.
-Notably, we limit the evaluation exhibited in the Experiment section to merely three metrics, including Accuracy, Feature Distance and FID (Heusel et al., 2017), while Sample Diversity is presented in the Sec.C.1 in the Appendix.
-Moreover, the VMI (Wang et al., 2021a) and RLBMI (Han et al., 2023) will be further evaluated in Sec.
-C.7 owing to their excessive need of time.
+Following the common setting in MI field, our benchmark concentrates on the evaluation at distribution level instead of sample level, assessing the overall performance of the whole reconstructed dataset.
+Therefore, we provide a total of 9 widely recognized distribution level evaluation metrics for users, which can be categorized into four types according to the evaluated content:
+Attack Accuracy.
+This metric serves as the primary measurement for MI attacks, using a given classifier to classify the inverted images and measures the top- () and top- () accuracy for target labels.
+The higher the reconstructed samples achieve attack accuracy on another classifier trained with the same dataset, the more private information in the dataset can be considered to be exposed [63].
+Therefore, higher attack accuracy denotes better attack performance.
+Feature Distance.
+The feature is defined as the output of the given classifier’s penultimate layer.
+We compute the shortest feature distance between inverted images and private samples for each class and calculate the average distance.
+In our benchmark, we follow the common choices across previous methods [43, 39] and select the evaluation model and another pre-trained FaceNet [40] to compute the feature distance, respectively denoted as and .
+Smaller feature distance means more similar features to the private image.
+Fréchet Inception Distance (FID).
+FID [19] is commonly used to evaluate the generation quality of GANs by computing the distance between feature vectors from all the target private data and that from all the reconstructed samples.
+Feature vectors are extracted by an Inception-v3 [46] pre-trained on ImageNet [7].
+Lower FID score shows higher inter-class diversity and realism [49].
+Sample Diversity.
+The metric contains Precision-Recall [27] and Density-Coverage [32] scores.
+Higher values indicate greater intra-class diversity of the inverted images.
+Since this metric serves as a secondary indicator for auxiliary reference in MI tasks, we have placed the corresponding evaluation results in the Supp.C.1.
+4 Evaluation
+4.1 Evaluation Setups
+To ensure fair and uniform comparison and evaluation, we select the FFHQ [22] as the public dataset and FaceScrub [33] as the private dataset for all the experiments in the Experiment section.
+The target models are fixed to the IR-152 [17] for low-resolution scenario and ResNet-152 [17] for high-resolution scenario, both trained on the FaceScrub.
+For each attack method, the number of images reconstructed per class is set to 5 due to time and computation cost.
+More detailed experimental settings are listed in Supp.B.
+Notably, we limit the evaluation exhibited in the Experiment section to merely three metrics, including Accuracy, Feature Distance, and FID [19], while Sample Diversity is presented in the Supp.C.1 due to the paper length limitation.
+To measure the deviation across the target labels, we introduce error bars for Accuracy and Feature Distance.
+As FID metric is computed based on all the recovered images, it is not essential to calculate error bars.
+Moreover, VMI [49] and RLBMI [16] will be further evaluated in Supp.
+C.6 owing to their excessive need of time (See Supp.
+C.5).
+We also conduct a user study in Supp.
+D.4.
 4.2 Evaluation on Different Attack Methods
-In this part, we prepare a unified experimental setting for different MI attack methods to conduct a fair comparison.
-The resolution of private and public datasets is set to , indicating a relatively easier scenario.
+In this part, we prepare a unified experimental setting for different MI attack methods to ensure fair comparison.
+The resolution of private and public datasets is set to , indicating relatively easier scenarios.
 Comparisons of white-box and black-box MI attacks are presented in Table 3.
-Remarkably, the PLGMI (Yuan et al., 2023a) and LOKT (Nguyen et al., 2023a) achieve state-of-the-art comprehensive performance in white-box attacks and black-box attacks respectively, showing significant superiority in Accuracy and Feature Distance metrics.
-However, the lowest FID scores occur in the PPA (Struppek et al., 2022) and C2FMI (Ye et al., 2023) respectively instead of the above methods.
+Remarkably, the PLGMI [60] and LOKT [34] achieve state-of-the-art comprehensive performance in white-box attacks and black-box attacks respectively, showing significant superiority in Accuracy and Feature Distance metrics.
+However, the lowest FID scores occur in the PPA [43] and C2FMI [59] respectively instead of the above methods.
 We infer that this is because PPA and C2FMI employ more powerful generators (e.g.
-StyleGAN2 (Karras et al., 2020b)) as the GAN prior compared to PLGMI and LOKT, leading to more real image generation.
-Visualization in Fig 2 further validates the inference.
+StyleGAN2 [24]) as the GAN prior compared to PLGMI and LOKT, leading to more real image generation.
+Visualization results displayed in Fig 2 further validate the inference.
+Notably, the visualization of Mirror(black) in Fig.2 and Fig.6 (in Supp.C.7) is anomalous compared to other methods.
+We perform further investigation on this special phenomenon in Supp.D.3.
+The experimental results demonstrate that this phenomenon is highly likely attributable to the genetic algorithm compromising the integrity of StyleGAN’s latent vectors, consequently inducing a deviation in image generation from the normal distribution.
 | Method | |||||
@@
 | IF-GMI | |||||
-| BREPMI | |||||
 | Mirror(black) | |||||
 | C2FMI | 43.488 | ||||
-| LOKT |
+| BREPMI | |||||
+| LOKT | 71.701 |
 4.3 Evaluation on Higher Resolution
@@
 All the methods that employ stronger GAN priors maintain low FID scores, including Mirror, C2FMI, and PPA, while other methods suffer from significant degradation in image reality.
-This phenomenon is more pronounced in the visualization results displayed in Fig 3.
 Despite the primary metric for evaluating MI attacks is Accuracy, the reality of reconstructed images should be ensured within a reasonable range for better image quality.
 Thus, it is imperative to explore more complex GAN priors with enhanced performance in future research, extending the MI field to more challenging and practical applications.
+More corresponding visualization results are displayed in Supp.C.7.
 | Method | |||||
@@
 | IF-GMI | |||||
-| BREPMI | |||||
 | Mirror (black) | |||||
 | C2FMI | 47.317 | ||||
+| BREPMI | 105.489 | ||||
 | LOKT |
-4.4 Evaluation on Different Model Predictive Power
+4.4 Evaluation on Varied Model Predictive Power
 The predictive power of the target model is a crucial factor in determining the effectiveness of MI attacks.
-Previous work (Zhang et al., 2020a) has conducted preliminary experimental validations on simple networks (e.g.
-LeNet (LeCun et al., 1989)), demonstrating that the performance of the first GAN-based attack GMI is influenced by the predictive power of the target model.
+Previous work [63] has conducted preliminary experimental validations on simple networks (e.g.
+LeNet [28]), demonstrating that the performance of the first GAN-based attack GMI is influenced by the predictive power of the target model.
 Therefore, we evaluate the state-of-the-art MI attacks on target models with varied predictive power to further validate the consistency of this characteristic in the recent attacks.
-The resolution of datasets is set to .
-The evaluation is stated in Fig.
-4.
+The resolution of datasets is set to and the evaluation results are shown in Fig.
+3.
 The comparison in Fig.
@@
 Thus one can expect lower privacy leakage in the robust model by balancing the accuracy-privacy trade-off.
-This study of the predictive power of target models illustrates another useful aspect of our MIBench.
 4.5 Evaluation on Defense Strategies
-This analysis concentrates on the robustness of MI attacks when applied to target models with defense strategies.
+This analysis focuses on the robustness of MI attacks when applied to models with defenses.
 Notably, we select the first 100 classes subset from FaceScrub as the target dataset due to the time cost.
-The assessment results are listed in Fig.5.
-The configuration of defenses is set following the official parameters, as detailed in Table 8.
-Overall, the TL (Ho et al., 2024) achieves the state-of-the-art decrease in Accuracy.
+The assessment results are listed in Fig.4 and Fig.5.
+The configuration of defenses is set following the official parameters, as detailed in Table 9 in the Supplementary.
+From Fig.4, the TL [20] achieves the state-of-the-art decrease in Accuracy.
 However, advanced attack methods have overcome current defense strategies to some extent, such as PLG and PPA.
 Additionally, some older defense strategies (e.g.
-MID (Wang et al., 2021c)) are no longer effective against the latest attacks.
-From Fig.5, we observe that LS (Struppek et al., 2023) exhibits unexpectedly poor performance in metric while it was recently published in the top-tier conference.
-The potential reason might be the utilized target classifier with relatively high test accuracy, as validated in the Sec.
-C.4.
+MID [52]) are no longer effective against the latest attacks.
+As for Table 5, the SSD [65] is especially effective when defending against black-box and label-only attacks.
+This indicates that employing defense in the inference stage is more effective than in the training process.
+From Fig.4, we observe that LS [44] exhibits unexpectedly poor performance in metric while it was recently published in the top-tier conference.
+The potential reason might be the utilized target classifier with relatively high test accuracy, as validated in the Supp.D.1.
 Furthermore, we conduct further experiments with PPA as the attack method against ResNet-152 trained on high resolution (224×224) scenario, proving that this phenomenon can be extended into other defenses.
-The results are listed in the Table 5, demonstrating that all the defenses become invalid even with the recommended parameters.
-More in-depth evaluation on defenses is exhibited in Sec C.5.
-Combining this phenomenon with the above experiment on predictive power, our empirical analysis indicates that the leaked information is strongly correlated to the model prediction accuracy and current defenses cannot effectively reduce the privacy information without sacrifice of model performance.
-Our findings emphasize that more reliable and stable defense strategies should be studied due to the fact that high model prediction accuracy is crucial for application of AI technology.
+The results are listed in the Table 5, demonstrating that all the train-time defenses become invalid even with the recommended parameters.
+More in-depth evaluation on defenses is exhibited in Supp.D.2.
+Combining this phenomenon with the above experiment on predictive power, our empirical analysis indicates that the leaked information is strongly correlated to the model prediction accuracy, and current defenses cannot effectively reduce the privacy information without sacrificing of model performance.
+Our findings emphasize that more reliable and stable defense strategies should be studied.
+A potential solution is designinga more robust model architecture instead of merely adding a regularizer in the training process.
+4.6 Evaluation on Adversarial Robustness
+Most defense methods are designed to focus solely on countering MI attacks, while neglecting protection against other types of attacks.
+However, since both training-time and inference-time defenses can influence the model’s outputs, it becomes imperative to evaluate the robustness of defense-protected models when confronted with adversarial attacks.
+This necessity highlights the importance of comprehensive security assessments to ensure that enhanced privacy protections do not inadvertently compromise the model’s resilience against diverse attacks.
+We select the typical adversarial attacks FGSM [14], PGD [31], BIM [26], and OnePixel [45] as representative to evaluate different MI defenses.
+The target models are set to IR- trained under multiple defenses on the high-resolution FaceScrub dataset.
+The results are shown in Table 6.
+In general, the inference-time defenses demonstrate extraordinary enhancement in adversarial robustness, while train-time defenses fail to improve adversarial robustness.
+Especially, most train-time defenses reduce adversarial robustness under the untargeted scenario of OnePixel.
+Our findings reveal that train-time defenses fail to adequately address multi-faceted robustness requirements.
+A promising solution is integrating adversarial training into the train-time defense process to enhance robustness.
 | Method | Hyperparameters | Test Acc | |||||
-| NO Defense | - | 98.510 | 0.972 | 0.990 | 307.714 | 0.588 | 50.259 |
-| MID | 96.760 | 1.000 | 1.000 | 273.687 | 0.517 | 49.239 | |
-| 95.680 | 0.990 | 1.000 | 276.889 | 0.526 | 51.227 | ||
-| BiDO | 98.030 | 0.986 | 0.996 | 306.827 | 0.554 | 50.943 | |
-| 97.430 | 0.968 | 0.996 | 320.191 | 0.594 | 50.132 | ||
-| TL | 97.620 | 0.994 | 1.000 | 282.394 | 0.513 | 53.023 | |
-| 97.530 | 0.982 | 0.996 | 306.528 | 0.561 | 51.489 |
+| NO Defense | - | 0.985 | 0.972 | 0.990 | 307.714 | 0.588 | 50.259 |
+| MID | 0.967 | 1.000 | 1.000 | 273.687 | 0.517 | 49.239 | |
+| 0.956 | 0.990 | 1.000 | 276.889 | 0.526 | 51.227 | ||
+| BiDO | 0.980 | 0.986 | 0.996 | 306.827 | 0.554 | 50.943 | |
+| 0.974 | 0.968 | 0.996 | 320.191 | 0.594 | 50.132 | ||
+| LS | 0.967 | 0.971 | 0.996 | 281.956 | 0.524 | 38.924 | |
+| 0.964 | 0.949 | 0.995 | 302.442 | 0.563 | 39.185 | ||
+| TL | 0.976 | 0.994 | 1.000 | 282.394 | 0.513 | 53.023 | |
+| 0.975 | 0.982 | 0.996 | 306.528 | 0.561 | 51.489 | ||
+| TTS | 0.959 | 0.861 | 0.967 | 143.637 | 0.706 | 41.243 | |
+| 0.968 | 0.857 | 0.968 | 147.672 | 0.707 | 42.002 |
+| Method | Test Acc | Untargeted Attacks | Targeted Attacks | ||||||
+| FGSM | PGD | BIM | OnePixel | FGSM | PGD | BIM | OnePixel | ||
+| No defense | 0.982 | 0.997 | 0.995 | 1.000 | 0.665 | 0.087 | 0.170 | 0.161 | 0.000 |
+| MID | 0.982 | 1.000 | 0.997 | 0.997 | 0.695 | 0.087 | 0.139 | 0.153 | 0.002 |
+| BiDO | 0.940 | 1.000 | 1.000 | 1.000 | 0.826 | 0.085 | 0.114 | 0.112 | 0.007 |
+| LS | 0.980 | 0.997 | 0.997 | 0.997 | 0.665 | 0.092 | 0.158 | 0.148 | 0.007 |
+| TL | 0.953 | 1.000 | 1.000 | 1.000 | 0.873 | 0.117 | 0.175 | 0.170 | 0.007 |
+| TTS | 0.973 | 0.826 | 0.985 | 0.992 | 0.100 | 0.034 | 0.497 | 0.502 | 0.000 |
+| Purifier | 0.946 | 0.939 | 0.897 | 0.934 | 0.075 | 0.002 | 0.060 | 0.068 | 0.000 |
+| SSD | 0.950 | 0.902 | 0.995 | 0.987 | 0.209 | 0.004 | 0.087 | 0.061 | 0.000 |
 5 Conclusion
@@
 To the best of our knowledge, this is the first benchmark and first open-source toolbox in the MI field.
-Our benchmark encompasses 16 of the state-of-the-art MI attack methods and defense strategies and more algorithms will be continually updated.
-Based on the implemented toolbox, we establish a consistent experimental environment and conducted extensive experimental analyses to facilitate fair comparison between different methods.
-In our experiments, we explore the impact of multiple settings, such as different image resolutions, model predictive power and defense.
-With in-depth analysis, we have identified new insights and proposed potential solutions to alleviate them.
+Our benchmark encompasses 19 of the state-of-the-art MI attack methods and defense strategies, and more algorithms will be continually updated.
+Based on the implemented toolbox, we establish a consistent experimental environment and conducted extensive evaluation to facilitate fair comparison between different methods.
+With in-depth analyses, we have benchmarked all the typical MI attacks and defenses under the unified setups and identified new insights for future study.
 Societal Impact and Ethical Considerations.
-A potential negative impact of our benchmark could be malicious users leveraging the implemented attack methods to reconstruct private data from public system.
-To alleviate this potential dilemma, a cautious approach for data users is to adopt robust and reliable defense strategies, as shown in the Sec.4.5 of our paper.
-Additionally, establishing access permissions and limiting the number of visits for each user is crucial to build responsible AI systems, thereby alleviating the potential contradictions with individual data subjects.
+A potential negative impact could be malicious users leveraging the implemented attack methods to reconstruct private data from the public system.
+To alleviate potential dilemma, a cautious approach is to establish access permissions.
 Appendix Table
@@
 A.1 Details of Data Processing
-This section introduces three types of dataset processing implemented by our toolbox, including data pre-processing, dataset splitting and dataset synthesis.
+This section introduces three types of dataset processing implemented by our toolbox, including data pre-processing, dataset splitting, and dataset synthesis.
 A.1.1 Data Pre-processing.
@@
 Here are the default pre-processing method for the four datasets.
-FaceScrub (Ng & Winkler, 2014). In the low-resolution version, the original data is center cropped at and then scaled, and the high-resolution is scaled directly.
+FaceScrub [33]. In the low-resolution version, the original data is center cropped at and then scaled, and the high-resolution is scaled directly.
+AFHQ-Dog [4]. The images will be resized to the corresponding resolution.
+Stanford Dogs [6]. We crop the images following the official annotations and resize them to the corresponding resolution.
 A.1.2 Dataset Splitting.
 This step is used to model training, including classifiers and conditional generators.
-For labeled datasets like CelebA and FaceScrub, we provide a fixed partition to slice the training and test sets.
+For labeled datasets like CelebA and FaceScrub, we use 80% data for training, 10% for validation, and 10% for testing.
+The validation set is used to select the best trained models, hyperparameters for training, and hyperparameters for defenses.
 For unlabeled datasets such as FFHQ and MetFaces, we provide a script for pseudo-labeling all images by computing the scores.
-The strategy of PLGMI (Yuan et al., 2023a) for the score calculation is contained in the examples of our toolbox.
+The strategy of PLGMI [60] for the score calculation is contained in the examples of our toolbox.
 A.1.3 Dataset Synthesis.
-Some methods use synthetic datasets for model training. We provide a script to synthesis datasets by pre-trained GANs according to LOKT (Nguyen et al., 2023a).
+Some methods use synthetic datasets for model training. We provide a script to synthesis datasets by pre-trained GANs according to LOKT [34].
 A.2 Implementation of Attack Models.
-We implemented and versions for all the custom attack models defined by the official code of implemented algorithms.
-For those algorithms that use StyleGAN2 (Karras et al., 2020b), we provide a wrapper for loading the official model and adapting it to our attack process.
+By modifying the architecture of the generators and discriminators following PPA [43], we implemented and versions for all the custom attack models defined by the official code of implemented algorithms.
+For those algorithms that use StyleGAN2 [24], we provide a wrapper for loading the official model and adapting it to our attack process.
 A.3 Implementation of Classifier Training & Defense Methods
@@
 2.
-MID (Wang et al., 2021c).
+MID [52].
 It applies a Gaussian perturbation to the features in front of the classification head to reduce the mutual information of the model inputs and outputs.
@@
 3.
-BiDO (Peng et al., 2022).
+BiDO [37].
 It adds the regular term loss (COCO or HSIC) function so that the intermediate features decrease the mutual information with the input and increase the mutual information with the output.
@@
 4.
-Label Smoothing (LS) (Struppek et al., 2023). It adds an label smoothing term with a negative smoothing factor into the cross-entropy loss.
+LS [44]. It adds a label smoothing term with a negative smoothing factor into the cross-entropy loss.
 5.
-Transfer Learning (TL) (Ho et al., 2024).
+TL [20].
 This method uses a pre-trained model for fine-tuning.
@@
 In our implementation, the hyperparameter is defined as a ratio of the number of frozen layers.
+6.
+TTS [25].
+TTS first train classifier without defense.
+After that, the residual shortcuts in the last stages will be removed.
+Then, the classifier will be continuously trained.
+The hyperparameter means the keep ratio of a residual shortcut, and denotes the number of residual shortcuts to be removed.
+7.
+Purifier [57].
+Despite the lack of open code and details about and NN, we reproduce it according to their paper.
+The denotes the weight of keeping the first label when training the CVAE.
+If the L2 distance between the output and the nearest members, we swap the first and second labels.
+The validation set is used as the reference set it needs.
+8.
+SSD [65]. It modifies outputs by solving convex optimization problems. is used for probability calibration and is the distortion bound. The validation set is used to estimate .
 In order to fairly compare the various defense methods, all defense methods are trained using the same pre-trained model in our experiments.
@@
 Latent Vectors Sampling.
-This step generate random latent vectors and distributes them for each attack target.
+This step generates random latent vectors and distributes them for each attack target.
 Most attack methods use a random distribution strategy.
-Mirror, PPA and BREP calculate the score of each latent vector corresponding to each label, and for each label the vectors with the highest scores are selected.
+Mirror, PPA, and BREP calculate the score of each latent vector corresponding to each label, and for each label,l the vectors with the highest scores are selected.
 Labeled Latent Vectors Selection (Optional).
 The previous step distributes latent vectors for each label, and this step optimizes the latent vectors by calculating the scores of the latent vectors corresponding to the labels and selecting the few vectors with the highest scores.
-Although currently there are no algorithms use this step, it can be added into the attack algorithms that use conditional generators, e.g.
-PLGMI (Yuan et al., 2023a) and LOKT (Nguyen et al., 2023a).
+Although currently there are no algorithms that use this step, it can be added into the attack algorithms that use conditional generators, e.g., PLGMI [60] and LOKT [34].
 Optimization.
-Optimization is the key step for the attack process, it accepts the initialized latent vectors and attack labels as input and outputs the optimized inverted images.
+Optimization is the key step for the attack process.
+It accepts the initialized latent vectors and attack labels as input and outputs the optimized inverted images.
 We provide several kinds of optimization strategies of each attack method as follows.
@@
 It optimizes results from a Gaussian distribution of latent vectors corresponding to the target labels, and images are generated by random sampling from the optimized distribution.
-It is used by KEDMI (Chen et al., 2021b).
+It is used by KEDMI [3].
 - 3.
@@
 5.
-BREP Optimization. A specific optimizer for BREPMI algorithm (Kahla et al., 2022). It uses a boundary repelling strategy for gradient simulation.
+BREP Optimization. A specific optimizer for BREPMI algorithm [21]. It uses a boundary repelling strategy for gradient simulation.
 6.
-Reinforcement Learning Optimization. Optimizing the latent vectors via reinforcement learning. Used by RLB attack method (Han et al., 2023).
+Reinforcement Learning Optimization. Optimizing the latent vectors via reinforcement learning. Used by RLB attack method [16].
 Final Images Selection (Optional).
-This step works by calculating the scores of each image and selecting the part of the image with the highest score as the result of the attack. It is used by PPA (Struppek et al., 2022).
+This step works by calculating the scores of each image and selecting the part of the image with the highest score as the result of the attack. It is used by PPA [43].
 | Method | Latent Optimizer | Identity Loss | Prior Loss | Image Augment |
@@
 | LOKT | Adam | Max Margin | - | ✔ |
-A.5 Details of Evaluation
-We provide the following four evaluation metrics to evaluate the effectiveness of the attack.
-1.
-Classification Accuracy.
-The metric uses a given classifier to classify the inverted images and measures the top- and top- accuracy for target labels.
-The higher the reconstructed samples achieve attack accuracy on another classifier trained with the same dataset, the more private information in the dataset can be considered to be exposed (Zhang et al., 2020a).
-2.
-Feature Distance.
-The feature is defined as the output of the given classifier’s penultimate layer.
-We compute the shortest feature distance between inverted images and private samples for each class and calculate the average distance.
-Smaller feature distance means more similar features to the private image.
-- 3.
-- 4.
-Appendix B Experimental Details
+Appendix B Evaluation Details
 This section describes the setup and details of the experiments in this paper.
-B.1 Experiment Settings
-We conducted experiments at both low and high resolution scenarios.
-For the low resolution experiments, we employed classifiers with a resolution of as the target models, and an ResNet- with a resolution of served as the evaluation model.
+B.1 Evaluation Settings
+Following the standard settings in existing works, we conducted experiments at both low and high resolution scenarios.
+For the low-resolution experiments, we employed classifiers with a resolution of as the target models, and a ResNet- with a resolution of served as the evaluation model.
 In the high resolution experiments, we used classifiers with a resolution of as the target models, with an Inception-v3 model having a resolution of as the evaluation model.
@@
 | Dataset | Model | Resolution | Test Acc |
-| CelebA | VGG-16 | 0.8826 | |
-| ResNet-152 | 0.9371 | ||
-| ResNet-50 | 0.9588 | ||
-| ResNet-152 | 0.9003 | ||
-| Inception-v3 | 0.9216 | ||
-| FaceScrub | VGG-16 | 0.8785 | |
-| ResNet-152 | 0.9825 | ||
-| ResNet-50 | 0.9938 | ||
-| ResNet-152 | 0.9225 | ||
-| ResNeSt-101 | 0.9329 | ||
-| Inception-v3 | 0.9445 |
+| CelebA | VGG-16 | 0.882 | |
+| ResNet-152 | 0.937 | ||
+| ResNet-50 | 0.958 | ||
+| ResNet-152 | 0.900 | ||
+| Inception-v3 | 0.921 | ||
+| FaceScrub | VGG-16 | 0.878 | |
+| ResNet-152 | 0.982 | ||
+| ResNet-50 | 0.993 | ||
+| ResNet-152 | 0.922 | ||
+| ResNeSt-101 | 0.932 | ||
+| Inception-v3 | 0.944 |
 | Defense Method | Hyperparameters | Test Acc |
@@
 | LS | ||
-| TL |
-B.4 Evaluation.
-The definitions of the evaluation metrics are detailed in Section A.5. Here, we present the specific details of the metrics used in the experiments.
+| TL | ||
+| TTS | ||
+| Purifier | ||
+| SSD |
+B.4 Evaluation Metrics
+The definitions of the evaluation metrics are detailed in Section 3.3. Here, we present the specific details of the metrics used in the experiments.
 For the Classification Accuracy and Feature Distance metrics, we evaluate the attack results using another classifier pre-trained on the same dataset as the target model: ResNet-50 for low-resolution settings and Inception-v3 for high-resolution settings, denoted as Acc and .
@@
 We use Inception-v3, pre-trained on ImageNet, to extract the features of images and participate in the score calculation.
-Appendix C More Experimental Results
+Appendix C More Evaluation
 C.1 Sample Diversity
-Following the settings of Section 4.2 and 4.3, we computed the Precision-Recall (Kynkäänniemi et al., 2019) and Density-Coverage (Kynkäänniemi et al., 2019) to evaluate the intra-class diversity for each attack method.
-The results are presented in Table 9, 10, 11 and 12.
+Following the settings of Section 4.2 and 4.3, we computed the Precision-Recall [27] and Density-Coverage [27] to evaluate the intra-class diversity for each attack method.
+The results are presented in Table 10, 11, 12 and 13.
 It tends to be that the attacks with stronger GAN priors get higher scores, such as Mirror, C2FMI and PPA.
@@
 In addition to attacking IR-152 and ResNet-152 in Section 4.2 and 4.3, we extend our experiments on more different target classifiers.
-We evaluate attacks on VGG- (Simonyan & Zisserman, 2014) and ResNeSt- (Zhang et al., 2022a) in low and high resolution settings, respectively.
-The results are presented in Table 13 and 14.
+We evaluate attacks on VGG- [41] and ResNeSt- [62] in low- and high-resolution settings, respectively.
+The results are presented in Table 14 and 15.
 | Method | |||||
@@
 Besides the aforementioned CNN-based classifiers, we further analyze MI attacks on ViT-based models.
-Table 15 presents further experiments on the MaxViT (Tu et al., 2022) under the high resolution (224×224) scenario.
+Table 16 presents further experiments on the MaxViT [48] under the high resolution (224×224) scenario.
 Other experimental settings are continuous with the main paper, with FFHQ as the public dataset and FaceScrub as the private dataset.
@@
 Therefore, we design more combination of datasets in this part to further assess the transferability of different attacks.
-The results are listed in Table 16 and 17.
-The visual results are shown in Figure 6 and 7.
-Except for the typical face classification task, we have conducted more experiments on the dog breed classification task under the high resolution (224×224) scenario, which includes two non-facial datasets, Stanford Dogs (Dataset, 2011) and Animal Faces-HQ Dog (AFHQ) (Choi et al., 2020).
+The results are listed in Table 17 and 18.
+The visual results are shown in Figure 9 and 10.
+Except for the typical face classification task, we have conducted more experiments on the dog breed classification task under the high resolution (224×224) scenario, which includes two non-facial datasets, Stanford Dogs [5] and Animal Faces-HQ Dog (AFHQ) [4].
 The public dataset is AFHQ while the private dataset is Stanford Dogs.
@@
 | LOMMA+KEDMI | ||||
-| IF-GMI |
-C.4 Evaluation on LS Defense Method
-To evaluate the LS defense method, we follow the settings in the official paper (Struppek et al., 2023) and evaluate it in more settings, shown in Table 19.
-To better evaluate the effectiveness of LS defense algorithms by making the undefended classifiers slightly less accurate than the defense-imposed classifiers through an early-stop strategy when training.
-In setting A, the public and private datasets are different part of CelebA dataset in low resolution scenario with the GMI as the attack method.
-In setting B, the public dataset is FFHQ and the private dataset is FaceScrub with the PPA as the attack method.
-In these two settings, the predictive power is much lower than that describes in Section 4.5.
-In this case, the LS defense is very effective, making the success rate of the attack drop dramatically.
-| Setting | Defense | Test Acc | ||||
-| A | - | |||||
-| LS | ||||||
-| B | - | |||||
-| LS |
-C.5 Validation for Defense Methods on Models with Low Accuracy
-The results for models with low accuracy are listed in Table 20. Obviously, the defense is valid when the target model has relatively low prediction accuracy.
-| Method | Hyperparameters | Test Acc | |||||
-| NO Defense | - | 92.170 | 0.686 | 0.914 | 262.471 | 0.767 | 68.454 |
-| MID | 88.240 | 0.568 | 0.820 | 246.021 | 0.757 | 69.663 | |
-| BiDO | 88.620 | 0.582 | 0.874 | 275.453 | 0.793 | 68.248 | |
-| TL | 89.160 | 0.316 | 0.616 | 279.439 | 0.897 | 63.241 |
-C.6 Evaluation on Different Loss Functions
+| IF-GMI | 48.972 |
+C.4 Evaluation on Different Loss Functions
 In recent years, various attack algorithms have attempted to mitigate the effects of gradient vanishing by employing different loss functions.
@@
 Specifically, we adopt PPA with FFHQ prior to attack a ResNet-152 classifier pre-trained on FaceScrub.
-The comparison of the results is presented in Table 21.
-Our findings indicate that the Poincare loss function yields the highest performance without model augmentation, whereas the Logit loss function achieves the best results with model augmentation.
+The comparison is presented in Table 20.
+Our findings indicate that the Poincaré loss function yields the highest performance without model augmentation, whereas the Logit loss function achieves the best results with model augmentation, serving as the optimal loss choices for future studies.
 | Loss Function | |||||
@@
 | Logit+ |
-C.7 More evaluation on VMI, RLBMI and PPA
+C.5 Evaluation on Time Overhead
+We conduct an extra evaluation on the time overhead among all the MI attacks.
+The results are recorded in the Table 21.
+The time complexities of most algorithms fall within the same order of magnitude, which is around 100 seconds per class.
+Notably, the time cost of VMI and RLBMI is extraordinarily higher than other attacks, indicating that both methods are hard to evaluate under the identical scenarios in the main evaluation and relatively unpractical in real-world settings.
+Therefore, we put the relevant experiments on the next section (Supp.C.6).
+| Attacks | |
+| GMI | 78 |
+| KEDMI | 73 |
+| VMI | 5910 |
+| Mirror(white-box) | 43 |
+| PPA | 124 |
+| PLGMI | 59 |
+| LOMMA+GMI | 133 |
+| LOMMA+KEDMA | 127 |
+| IF-GMI | 83 |
+| Mirror(black-box) | 161 |
+| C2FMI | 83 |
+| RLBMI | 10633 |
+| BREPMI | 298 |
+| LOKT | 219 |
+C.6 More evaluation on VMI, RLBMI and PPA
 Considering the high computational overhead of RLBMI and VMI, we only experiment at a low resolution settings for classes.
-The public and private datasets are different part of CelebA dataset.
+The public and private datasets are splitted part of CelebA dataset.
 The results are shown in Table 22.
@@
 | 20 | 600 |
-Appendix D Limitations and Future Plans
-Our benchmark mainly focuses on GAN-based MI attacks and MI defenses applied in the classifier training stage.
-We will extend our benchmark to wider range of MI methods, including learning-based MI attacks and MI defenses applied to the classifier output.
-Furthermore, the concept of MI also pervades across modalities (e.g.
-text (Parikh et al., 2022; Zhang et al., 2023; Carlini et al., 2019; 2021; Yu et al., 2023; Nasr et al., 2023) and graph learning (Zhang et al., 2021; 2022b; Zhou et al., 2023; Wu et al., 2022; He et al., 2021)) beyond computer vision domain, where our benchmark concentrates in current stage.
-Expansion to new modalities is a promising direction for our benchmark to further explore the privacy threats in other fields, leading to more generalization in the AI security.
+C.7 More Visualization
+We visualize more results in Figure 6, 9 and 10. The target model is ResNet- trained on CelebA dataset in high resolution setting.
+Appendix D Special Experiments
+D.1 Validation on LS Defense Method
+To validate the correctness of our implemented LS defense, we follow the settings in the official paper [44] and evaluate it in more settings, shown in Table 24.
+To better evaluate the effectiveness of LS defense algorithms, we make the undefended classifiers slightly less accurate than the defense-imposed classifiers through an early-stop strategy when training.
+In setting A, the public and private datasets are different part of CelebA dataset in low resolution scenario with the GMI as the attack method.
+In setting B, the public dataset is FFHQ and the private dataset is FaceScrub with the PPA as the attack method.
+In these two settings, the predictive power is much lower than that describes in Section 4.5.
+In this case, the LS defense is very effective, making the success rate of the attack drop dramatically.
+| Setting | Defense | Test Acc | ||||
+| A | - | |||||
+| LS | ||||||
+| B | - | |||||
+| LS |
+D.2 Validation for Defense Methods on Models with Low Accuracy
+The results for models with low accuracy are listed in Table 25. Obviously, the defense is valid when the target model has relatively low prediction accuracy.
+| Method | Hyperparameters | Test Acc | |||||
+| NO Defense | - | 92.170 | 0.686 | 0.914 | 262.471 | 0.767 | 68.454 |
+| MID | 88.240 | 0.568 | 0.820 | 246.021 | 0.757 | 69.663 | |
+| BiDO | 88.620 | 0.582 | 0.874 | 275.453 | 0.793 | 68.248 | |
+| TL | 89.160 | 0.316 | 0.616 | 279.439 | 0.897 | 63.241 |
+D.3 The abnormal visualization of Mirror
+To determine the cause of visualization anomalies in the Mirror method, we conduct a simple ablation experiment on its optimization objective, as denoted in Table 26 and Fig 8.
+The experimental results demonstrate that when substituting the optimization objective with the space latent vectors, the visualization of Mirror returns to normality.
+Therefore, the underlying reason of the visualization anomalies is likely attributed to genetic-algorithm-induced corruption of StyleGAN’s latent vectors.
+As a disentangled latent vector that deeply participates in StyleGAN’s hierarchical progressive image generation process, the compromised latent vector ultimately leads to image artifacts in the generated outputs.
+| Scenario | Space | ||||
+| White-box | |||||
+| Black-box | |||||
+D.4 User Study
+In this section, we conduct a user study considering a subjective evaluation of MI attacks to further evaluate their effectiveness.
+We use different attack results in the resolution scenarios where the public dataset is FFHQ and the private dataset is FaceScrub.
+We randomly select classes for user study and publish survey questionnaires on the credamo website.
+The user interface is shown in Fig.
+7.
+For each class, users are shown private images (above in Fig.7) as examples and select the reconstructed images (below in Fig.7) that have the top-3 visual similarity to the examples.
+In total, users have participated in the annotation, and the result is shown in Table 27.
+| Attacks | |
+| GMI | 0.035 |
+| KEDMI | 0.084 |
+| Mirror(white-box) | 0.493 |
+| PPA | 0.564 |
+| PLGMI | 0.326 |
+| LOMMA+GMI | 0.074 |
+| LOMMA+KEDMA | 0.083 |
+| IF-GMI | 0.419 |
+| BREPMI | 0.026 |
+| Mirror(black-box) | 0.027 |
+| C2FMI | 0.066 |
+| LOKT | 0.059 |
+Appendix E Limitations and Future Plans.
+During the writing process, there were new works about MI published in the top-tier conferences, such as [38].
+As a toolbox for MI research, our benchmark is continuously updated and all the new works are scheduled to be implemented in the near future.
 In addition to developing new algorithms, it is also essential to conduct further research on the MI attacks and defenses to make in-depth analysis about their characteristics and bring valuable new insights.